### v2

# Continuous Audio Language Models

Abstract
Audio Language Models (ALM) have emerged as the dominant paradigm for speech and music generation by representing audio as sequences of discrete tokens. Yet, unlike text tokens, which are invertible, audio tokens are extracted from lossy codecs with a limited bitrate. As a consequence, increasing audio quality requires generating more tokens, which imposes a trade-off between fidelity and computational cost. We address this issue by studying Continuous Audio Language Models (CALM). These models instantiate a large Transformer backbone that produces a contextual embedding at every timestep. This sequential information then conditions an MLP that generates the next continuous frame of an audio VAE through consistency modeling. By avoiding lossy compression, CALM achieves higher quality at lower computational cost than their discrete counterpart. Experiments on speech and music demonstrate improved efficiency and fidelity over state-of-the-art discrete audio language models, facilitating lightweight, high-quality audio generation. Samples are available at hf.co/spaces/kyutai/calm-samples.
1 Introduction
Using classification over a finite vocabulary as the training objective for autoregressive sequence models is an effective approach for naturally discrete modalities such as text, where large-scale Transformer-based (Vaswani et al., 2017) language models such as LLaMa (Touvron et al., 2023) and GPT-4 OpenAI (2024) have achieved impressive results. To extend this powerful framework to continuous domains such as image, audio, or video, previous work has mostly relied on discretizing signals using lossy compression algorithms (van den Oord et al., 2018), such that they become akin to text. In particular, neural audio codecs (Zeghidour et al., 2021; Défossez et al., 2024a) have provided discrete representations of audio that are compact enough to allow for high-quality speech (Borsos et al., 2023; Wang et al., 2023) and music (Agostinelli et al., 2023; Copet et al., 2023) generation with autoregressive models. In this context, a Residual Vector Quantizer (RVQ) (Zeghidour et al., 2021) transforms an audio frame into a coarse-to-fine hierarchy of tokens. As quantization inevitably introduces a perceptual quality loss, generating high-fidelity audio requires increasing the bitrate of audio tokens, which amounts to using deeper hierarchies of RVQ tokens. A consequence of growing the size of the token matrix (along time and token depth) is an additional computational load for the generative model, as the strong dependencies between tokens of the same frame(Lemercier et al., 2024) prevent fully parallel generation. The naive approach of flattening the token hierarchy (Borsos et al., 2023) being prohibitively expensive,  Copet et al. (2023) introduces a delay pattern that conjugates the computational efficiency of parallel generation with a better modeling of inter-token dependencies. Lee et al. (2022) and Yang et al. (2023) furthermore introduce a smaller RQ-Transformer model that is autoregressive along the depth axis, and  Défossez et al. (2024b) combines this approach with the delay pattern. While these methods currently power state-of-the-art generative models for audio (Défossez et al., 2024b; Labiausse et al., 2025), the trade-off imposed by residual quantization between quality and computation remains too constraining for generating high-quality audio on edge devices.
This motivates an alternative strategy: autoregressive modeling of continuous latents without quantization. Standard variational autoencoders (VAEs) are easier to train, are not affected by issues such as codebook collapse, and can reconstruct audio at higher fidelity for the same latent dimensionality. Pioneering work in the vision domain that autoregressively models continuous sequences includes GIVT (Tschannen et al., 2024) and MAR (Li et al., 2024), followed by some larger models in the image domain (Fan et al., 2025; Gu et al., 2025) and attempts in the audio domain (Turetzky et al., 2024; Pasini et al., 2024b; Jia et al., 2025). In MAR, the authors model the per-token probability distribution with a diffusion model (in the form of a small MLP) conditioned on a latent variable modeled by an autoregressive transformer backbone.
SALAD (Turetzky et al., 2024) and DiTAR (Jia et al., 2025) adapt MAR-style diffusion heads for text-to-speech modeling, achieving better audio quality than discrete baselines. However, these works are limited to text-to-speech on small-scale and domain-specific datasets, leaving open the question of how well they can adapt to richer audio domains such as music.
In this work, we propose Continuous Audio Language Models (CALM) that predict sequences in the latent space of a VAE, bypassing the need for quantization. While we build on the MAR architecture where a transformer backbone uses $(\mathbf{x}^{1},\ldots,\mathbf{x}^{s-1})$ to predict an intermediate latent $\mathbf{z}^{s}$, which then conditions a head (diffusion model) that models $p(\mathbf{x}^{s}|\mathbf{z}^{s})$, we find that without further improvements, it fails to generate rich audio content. To overcome this, we introduce several key innovations:
1. Improving quality and stability: To mitigate error accumulation during inference, we follow Pasini et al. (2024b) and, during training, inject noise into the long-term context $(\mathbf{x}^{1},\ldots,\mathbf{x}^{s-1})$. Additionally, we introduce a short-context transformer that summarizes recent clean latents, providing the sampling head with both coarse long-range context and fine-grained local information. 2. Diffusion-to-Consistency replacement: We replace the diffusion model with a continuous consistency model (Lu & Song, 2025) during training, significantly accelerating inference without compromising sample quality. This change reduces the inference time of the sampler head by a factor of up to $\times 20$ in our music experiments and $\times 12$ in our speech experiments compared to the use of a model that uses an RQ-Transformer as head. This method is modality-agnostic and may benefit other domains such as image generation. 3. Gaussian Temperature sampling: Temperature control is crucial for high-quality speech generation, yet consistency models lack a formal mechanism for temperature sampling. We present a heuristic that approximates temperature sampling in the consistency setup. 4. Head batch multiplier: Sampling multiple noise levels at training time for the same latent highly accelerates training for a small cost.
Our model generates high-quality speech and music with lower inference time, outperforming state-of-the-art discrete-token autoregressive models in both efficiency and audio fidelity. To our knowledge, this is the first application of autoregressive continuous latent modeling to both domains.
Figure 1: Overview of our model. During training, latent vectors $\mathbf{x}^{s}$ are noised to encourage the backbone Transformer to focus on coarse structure. The consistency head is a consistency model conditioned on the latent variable $\mathbf{z_{\text{long}}^{s}}$ produced by the backbone, as well as a short-term context vector $\mathbf{z_{\text{short}}^{s}}$ computed from a short-context Transformer applied to the most recent clean latent tokens.
2 Related Work
Autoregressive audio language models. Early autoregressive audio models operated on raw waveforms such as WaveNet (van den Oord et al., 2016) or learned discrete codes via VQ-VAE, such as in Jukebox (Dhariwal et al., 2020). The advent of neural audio codecs (Zeghidour et al., 2021; Défossez et al., 2024a) enabled high-quality vector-quantized tokenizers for general audio. These, in turn, powered discrete-token audio LMs: AudioLM (Borsos et al., 2023) for unconditioned audio generation while AudioGen (Kreuk et al., 2023), MusicLM (Agostinelli et al., 2023), and MusicGen (Copet et al., 2023) apply similar methods for text-to-audio and text-to-music generation. In speech, autoregressive modeling of RVQ codes has been used for text-to-speech generation (Wang et al., 2023; Kharitonov et al., 2023) as well as for spoken dialogue (Défossez et al., 2024b) or translation (Labiausse et al., 2025). However, all of these systems rely on lossy quantization, which inevitably degrades audio fidelity unless a significant compute budget is spent to generate a deep hierarchy of RVQ tokens. This is unlike CALM, which predicts continuous embeddings in one pass, providing a better quality-computation trade-off.
Continuous autoregressive models. In GIVT, Tschannen et al. (2024) use a transformer to autoregressively model the latent space of a VAE trained on images, parameterize a Gaussian Mixture Model, and train it with a cross-entropy loss. The authors of MAR (Li et al., 2024) obtain better results by replacing the Gaussian mixture model with a diffusion-model enabling the approximation of more flexible distributions. In MAR, a large transformer backbone predicts a continuous embedding $\mathbf{z}^{s}$ given $(\mathbf{x}^{1},\ldots,\mathbf{x}^{s-1})$, which then conditions a small MLP diffusion network that models the probability distribution $p(\mathbf{x}^{s}|\mathbf{z}^{s})$ of the next latent. This eliminates the need for discrete tokenizers, but at the cost of slow sampling: MAR typically needs hundreds of denoising steps per token. Hang et al. (2025) aims to speed up MAR by replacing the diffusion head with a shortcut head (Frans et al., 2025) for few-step sampling. Shortcut models combine a diffusion loss and a self-consistency loss in order to accelerate the diffusion process. They reduce the number of diffusion steps from 100 to 8 with a similar image quality. Apart from this work, we are not aware of other works that try to reduce the inference time of the diffusion process in continuous autoregressive models. Remarkably, CALM achieves a quality comparable to the best discrete models with only one step of consistency modeling.
In audio, the approach of MAR has been adapted for the task of Text-to-Speech (TTS) synthesis. SALAD (Turetzky et al., 2024) introduces a zero-shot TTS model that operates on continuous speech representations using a per-token latent diffusion process. By leveraging semantic tokens for contextual information and determining synthesis stopping points, SALAD achieves improved intelligibility and audio quality without relying on quantization. Similarly, DiTAR (Jia et al., 2025) presents a patch-based autoregressive framework combining a language model with a diffusion transformer for speech generation. This approach models dependencies between aggregated local patches of continuous tokens, using a causal language model to produce embeddings, which, along with previous patches, serve as inputs to a bidirectional diffusion transformer that predicts the next patch. The authors observe that providing local context through patching was determinant to improve their model, which corroborates what we observe with the introduction of a short context transformer into our model. While these two studies focus on TTS modeling, we extend the method to music. In Pasini et al. (2024b), the authors apply the MAR framework to music generation using a relatively small dataset comprising 20,000 single-instrument stems, training their model on 10-second excerpts on top of their continuous compression model Music2Latent (Pasini et al., 2024a). They introduce a method for noise augmentation of the data that allows the model to avoid error accumulation. However, we notice that scaling to a more complex and diverse dataset consisting of full musical pieces makes their approach struggle to maintain high-quality generation on longer sequences. To address these challenges, we propose novel strategies that enhance generation quality while also improving inference efficiency. Finally, Music2Latent2 (Pasini et al., 2025) explores combining autoregressive and consistency modeling but for compression.
3 Background
Notations: Let $W\in\mathbb{R}^{f_{s}\cdot d}$ be a monophonic waveform of $d$ seconds sampled at frame rate $f_{s}$. Our goal is to model $W$ either in the discrete latent space of a RVQ-based codec or in the latent space of a VAE. Let $f_{r}$ be the frame rate of the codec or VAE.
In the case of the discrete modeling, $W$ is represented by the sequence of discrete tokens $\left(q^{s,k}\right)$, where $s\in\{1,\ldots,S\}$ indexes time and $k\in\{1,\ldots,K\}$ indexes codebook depth and $S=f_{r}\cdot d$. Each token $q^{s,k}\in\{1,\ldots,N_{k}\}$ is drawn from a finite vocabulary.
In the case of continuous modeling, $W$ is represented by a sequence $\left(\mathbf{x}^{1},\ldots,\mathbf{x}^{S}\right)$ with $S=f_{r}\cdot d$ and $\mathbf{x}^{s}\in\mathbb{R}^{C}$ where $C$ is the latent dimension of the VAE.
3.1 Autoregressive Modeling with Residual Vector Quantization based codecs.
Autoregressive modeling of discrete tokens from RVQ-based codecs (Zeghidour et al., 2021; Défossez et al., 2024a) is a prevalent method for high-fidelity audio generation (Copet et al., 2023; Agostinelli et al., 2023; Borsos et al., 2023; Kreuk et al., 2023). Given a sequence of discrete tokens $\left(q^{s,k}\right)_{i\in\{1,\ldots,N\},k\in\{1,\ldots,K\}}$, early models like Borsos et al. (2023); Agostinelli et al. (2023) flattened the multi-level sequence, increasing its length by a factor of $K$ and resulting in high computational costs due to the quadratic complexity of Transformer self-attention. MusicGen (Copet et al., 2023) mitigates this by using a delay pattern to independently sample each of the $K$ RVQ levels, adding only K1K-1 tokens to the sequence but introducing a fixed latency of K1K-1 frames, which is problematic for real-time applications. RQ-Transformer (Lee et al., 2022) addresses this by using a sampler transformer head that models the RVQ at a given time step, enabling low-latency generation.
Denoting $\mathbf{q}^{s}=(q^{s,1},\ldots,q^{s,K})$, the Backbone Transformer $T_{\theta}$ encodes the history of previous timesteps to produce a context vector $\mathbf{z}^{s}$, and a RQ-Transformer $g_{\phi}$ autoregressively decodes the residual-wise components of the token stack at each time step. The context vector is then given by $\mathbf{z}^{s}=T_{\theta}(\mathbf{q}^{1},\ldots,\mathbf{q}^{s-1})$, and the logits $\ell^{s,k}$ for predicting the $k$-th codebook token are computed as $\ell^{s,1}=\text{Lin}(\mathbf{z}^{s})$ and $\ell^{s,k}=g_{\phi}\left(\mathbf{z}^{s},q^{s,1},\ldots,q^{s,k-1}\right)$ for $k>1$.
These logits are trained using a cross-entropy loss over discrete tokens $\mathcal{L}_{\text{CE}}=-\sum_{s=1}^{S}\sum_{k=1}^{K}\log p\left(q^{s,k}\mid\mathbf{q}^{<s},\mathbf{q}^{s,<k}\right).$
This approach enables efficient parallel modeling of RVQ sequences, allowing all codebooks corresponding to a single timestep to be generated simultaneously without introducing any delay. However, a key limitation lies in the use of the RQ-Transformer, which is computationally intensive; as its resource requirements scale with the number of RVQ.
3.2 Consistency Models
Flow Matching and Probability Flow ODE.
Let $p_{\text{data}}$ be a data distribution over $\mathbb{R}^{d}$. Given $\mathbf{x}_{0}\sim p_{\text{data}}$, diffusion (Ho et al., 2020; Song et al., 2021) and flow matching (Lipman et al., 2023) models define a forward noising process that gradually perturbs samples from $p_{\text{data}}$ through the noising process $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ and $t\in[0,T]$. $\alpha_{t},\sigma_{t}$ are predefined functions such that $\alpha$ is decreasing with $\alpha_{0}=1,\alpha_{T}=0$ and $\sigma$ is increasing with $\sigma_{0}=0,\sigma_{T}=1$. In Flow Matching, a neural network $F_{\phi}$ is trained to minimize the loss $\mathcal{L}_{\text{FM}}(\phi)=\mathbb{E}_{\mathbf{x}_{0}\sim p_{\text{data}},\,\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),\,t\sim\mathcal{U}(0,1)}\left[w(t)\left\|F_{\phi}(\mathbf{x}_{t},t)-\left(\alpha^{\prime}_{t}\mathbf{x}_{0}+\sigma^{\prime}_{t}\bm{\epsilon}\right)\right\|_{2}^{2}\right].$
Once trained, sample generation is performed by solving a deterministic ordinary differential equation known as the probability flow ODE (PF-ODE), which defines a continuous path from noise to data. In the context of Flow Matching, the PF-ODE is $\frac{d\mathbf{x}_{t}}{dt}=F_{\phi}(\mathbf{x}_{t},t)$ with $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$.
A Continuous-Time Consistency Models. (Song et al., 2023) is a neural network $f_{\phi}(\mathbf{x}_{t},t)$ trained to map a noisy input $\mathbf{x}_{t}$ directly to the corresponding clean data $\mathbf{x}_{0}$ in a single step, by approximating the sampling trajectory of the probability flow ODE (PF-ODE) starting from $\mathbf{x}_{t}$. To ensure correct behavior, $f_{\phi}$ must satisfy the boundary condition $f_{\phi}(\mathbf{x},0)=\mathbf{x}$, thus leading to the common parameterization $f_{\phi}(\mathbf{x}_{t},t)=c_{\text{skip}}(t)\mathbf{x}_{t}+c_{\text{out}}(t)F_{\phi}(\mathbf{x}_{t},t)$, where $F_{\phi}$ is a neural network and the coefficients satisfy $c_{\text{skip}}(0)=1$ and $c_{\text{out}}(0)=0$ to fulfill the boundary condition.
By using $T=\frac{\pi}{2}$ and $\alpha_{t}=\cos(t),\sigma_{t}=\sin(t)$, Lu & Song (2025) derive the following continuous-time consistency loss where $w_{\psi}(t)$ is an adaptive weighting function:
$\mathcal{L}_{\text{CM}}(\phi,\psi)=\mathbb{E}_{\mathbf{x}_{t},t}\left[\!\frac{e^{w_{\psi}(t)}}{D}\left\|F_{\phi}\left(\mathbf{x}_{t},t\right)-F_{\phi^{-}}\left(\mathbf{x}_{t},t\right)-\cos(t)\frac{df_{\phi^{-}}(\mathbf{x}_{t},t)}{dt}\right\|_{2}^{2}-w_{\psi}(t)\right].$ (1)
3.3 Autoregressive Modeling of Continuous Latents via Diffusion.
Li et al. (2024) propose MAR, a method for autoregressive modeling over a sequence $\left(\mathbf{x}^{1},\ldots,\mathbf{x}^{S}\right)$ of continuous latent vectors extracted from a pretrained VAE, thereby eliminating the need for discrete quantization. As in the discrete case, a Backbone Transformer $T_{\theta}$ maps the context to an embedding: $\mathbf{z}^{s}=T_{\theta}(\mathbf{x}^{1},\ldots,\mathbf{x}^{s-1})$.
Then, a diffusion process parameterized by a neural network $\bm{\epsilon}_{\phi}$ is trained on each $\mathbf{x}^{s}$ with the loss $\mathcal{L}_{\text{diff}}(\theta,\phi)=\sum_{s=1}^{S}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(0,I),t\sim[0,1]}\left[\left\|\bm{\epsilon}-\bm{\epsilon}_{\phi}(\mathbf{x}_{t}^{s},\mathbf{z}^{s},t)\right\|^{2}\right]$ where $\mathbf{x}^{s}_{t}$ is a noisy version of $\mathbf{x}^{s}$ at diffusion timestep $t$: $\mathbf{x}_{t}^{s}=\alpha_{t}\mathbf{x}^{s}+\sigma_{t}\bm{\epsilon}$ with $\bm{\epsilon}\sim\mathcal{N}(0,I)$ and $\alpha_{t}$ and $\sigma_{t}$ are predefined schedules for all $t\in[0,1]$. In practice, an MLP significantly smaller than the backbone transformer estimates $\bm{\epsilon}_{\phi}$. This replaces the categorical prediction used in discrete models (done by the RQ-Transformer) with a denoising task in the continuous domain (done by the MLP). This method enables flexible and differentiable modeling of continuous signals without requiring to perform quantization on the latent space which can lead to several issues such as codebook collapse, balancing quantization losses and training instabilities. A key limitation of this approach is that sample quality depends on the number of diffusion steps at inference, raising the question of whether it can surpass the RQ-Transformer under similar computational constraints.
4 Method
4.1 Our VAE-GAN
Most autoregressive audio models are built upon RVQ-GAN architectures (Zeghidour et al., 2021; Défossez et al., 2024a; Kumar et al., 2023; Guo et al., 2025). Following the approach of Evans et al. (2024), we instead adopt a VAE-GAN framework, replacing the RVQ bottleneck with a VAE bottleneck to regularize the latent space and enforce a Gaussian prior. Our VAE is fully causal and draws from the architecture of Mimi (Défossez et al., 2024b), using Transformers in addition to convolutions in the encoder and decoder, which have been shown to improve performance.
While training the model with adversarial losses and VAE regularization without any reconstruction losses improves the quality of the model for speech, it degrades the reconstruction quality for music. Semantic distillation is performed for the speech VAE similarly as in Mimi, with WavLM (Chen et al., 2021) as teacher. There is no semantic distillation for the music model and we let this for future work as semantic content is harder to define for music. The loss is:
$\mathcal{L}_{\text{VAE}}=\lambda_{\text{t}}\mathcal{L}_{\text{t}}(x,\hat{x})+\lambda_{\text{f}}\mathcal{L}_{\text{f}}(x,\hat{x})+\lambda_{\text{adv}}\mathcal{L}_{\text{adv}}(\hat{x})+\lambda_{\text{feat}}\mathcal{L}_{\text{feat}}(x,\hat{x})+\lambda_{\text{KL}}\mathcal{L}_{\text{KL}}$ (2)
where $\mathcal{L}_{\text{t}}$ and $\mathcal{L}_{\text{f}}$ are the temporal and frequential reconstruction losses, $\mathcal{L}_{\text{adv}}$ is the adversarial loss, $\mathcal{L}_{\text{feat}}$ is the feature matching loss and $\mathcal{L}_{\text{KL}}$ is the KL regularization applied to the VAE bottleneck.
4.2 Consistency-Based Continuous Audio Language Model (CALM)
Let $\bigl{(}\mathbf{x}^{1},\ldots,\mathbf{x}^{S}\bigr{)}$ denote the sequence of continuous latent vectors produced by a VAE encoder. As illustrated in Fig. 1, our model comprises three main components:
1. Causal Backbone Transformer with Noise Injection
We build on the MAR framework (Li et al., 2024) by employing a causal Transformer $T_{\text{long},\theta^{1}}$ to capture long-term dependencies. During training, each input latent $\mathbf{x}^{s}$ is replaced with a noised version $\tilde{\mathbf{x}}^{s}$ (see Sec. 4.3 and Ablation 5). We don’t perform any noise augmentation at inference time. Concretely, $\mathbf{z}_{\text{long}}^{s}\;=\;T_{\text{long},\theta^{1}}\bigl{(}\tilde{\mathbf{x}}^{1},\,\ldots,\,\tilde{\mathbf{x}}^{s-1}\bigr{)}$. Noise injection prevents error accumulation during inference, but as shown in Tab. 5 is insufficient alone for high-quality music generation.
2. Short-Context Transformer
To supply local, high-resolution context to the denoising head, we introduce a lightweight causal Transformer that attends the $K$ previous clean latents (we use $K=10$, $\sim$ 0.4s of music): $\mathbf{z}_{\text{short}}^{s}\;=\;T_{\text{short},\theta^{2}}\bigl{(}\mathbf{x}^{s-K},\,\ldots,\,\mathbf{x}^{s-1}\bigr{)}$. This short-context embedding $\mathbf{z}_{\text{short}}^{s}$ supplies fine-grained information potentially lost through noise injection in the backbone. We show in Tab. 7 that the value of K is not a decisive hyperparameter.
3. Consistency-Model Head
Finally, a small MLP-based consistency model $f_{\phi}$ is conditioned on the sum of the long-term and short-term features, $\mathbf{Z}^{s}=\mathbf{z}_{\text{long}}^{s}+\mathbf{z}_{\text{short}}^{s}$. At inference time, for 1-step, the next latent $\hat{\mathbf{x}}^{s}$ is sampled through: $\bm{\epsilon}\sim\mathcal{N}(0,I),t=1$, $\hat{\mathbf{x}}^{s}\;=\;f_{\phi}\bigl{(}\mathbf{x}^{s}_{1}=\bm{\epsilon},t=1,\mathbf{Z}^{s}\bigr{)}$.
In addition to consistency, we experiment with the TrigFlow (Lu & Song, 2025) formulation of flow-matching for the MLP. Although TrigFlow yields marginally higher fidelity, its inference cost makes it impractical for real-time use. This tradeoff is studied in Sec. 5.2, Tab. 4.
Together, these three components form a continuous autoregressive model that (i) leverages noise-robust long-term modeling, (ii) preserves local detail via short-context conditioning, and (iii) achieves rapid, high-fidelity latent sampling through consistency modeling.
The training objective for one sequence $\bigl{(}\mathbf{x}^{1},\ldots,\mathbf{x}^{S}\bigr{)}$ is defined by:
$\small\mathcal{L}_{\text{CALM}}(\theta,\phi,\psi)=\sum_{s=1}^{S}\mathbb{E}_{t,\bm{\epsilon}}\left[\!\frac{e^{w_{\psi}(t)}}{D}\left\|F_{\phi}\left(\mathbf{x}^{s}_{t},t,\mathbf{Z}^{s}\right)-F_{\bar{\phi}}\left(\mathbf{x}^{s}_{t},t,\mathbf{Z}^{s}\right)-\cos(t)\frac{df_{\bar{\phi}}(\mathbf{x}^{s}_{t},t,\mathbf{Z}^{s})}{dt}\right\|_{2}^{2}-w_{\psi}(t)\right],$ (3)
where $\mathbf{Z}^{s}=\mathbf{z}_{\text{long}}^{s}+\mathbf{z}_{\text{short}}^{s}=T_{\text{long},\theta^{1}}\bigl{(}\tilde{\mathbf{x}}^{1},\,\ldots,\,\tilde{\mathbf{x}}^{s-1}\bigr{)}+T_{\text{short},\theta^{2}}\bigl{(}\mathbf{x}^{s-K},\,\ldots,\,\mathbf{x}^{s-1}\bigr{)}$,
$t\sim[0,1],\bm{\epsilon}\sim\mathcal{N}(0,I)$ and $\mathbf{x}_{t}^{s}=\cos(t)\mathbf{x}^{s}+\sin(t)\bm{\epsilon}$. All the parameters $(\theta,\phi,\psi)$ of the transformer backbone $T_{\text{long},\theta^{1}}$, the short-context transformer $T_{\text{short},\theta^{2}}$, the consistency MLP $f_{\phi}$ and the adaptive weighting function $w_{\psi}$ are jointly trained together with this consistency loss similarly as the backbone and the RQ-Transformer are trained through cross-entropy loss in the discrete case.
4.3 Combining noisy long-term context and clean short-term context
In preliminary experiments, we realized that the music generation model was diverging quite quickly during inference because it was not robust to error accumulation. Pasini et al. (2024b) introduce a noise augmentation trick at training time in order to tackle this problem. Given a sequence $(\mathbf{x}^{1},\mathbf{x}^{2},\ldots,\mathbf{x}^{S})$, they sample $k_{s}\sim\mathcal{U}(0,1)$ and $\bm{\epsilon}_{s}\sim\mathcal{N}(0,\mathbf{I})$ for every $s\in\{1,\ldots,S\}$ and compute $\tilde{\mathbf{x}}^{s}=k_{s}\bm{\epsilon}_{s}+(1-k_{s})\mathbf{x}^{s}$ for every $s$. After comparison and to preserve the variance of the $\mathbf{x}^{s}$ that are gaussians, we modify the noising formula as follows: $\tilde{\mathbf{x}}^{s}=\sqrt{k_{s}}\bm{\epsilon}_{s}+\sqrt{1-k_{s}}\mathbf{x}^{s}$
In our case, applying this noising method to music generation slightly improves model stability but often reduces detail and instrument diversity, typically preserving only rhythmic elements, with audio fading into silence after 10–15 seconds. Since Pasini et al. (2024b) targets short, single-instrument clips, it’s unsurprising the method performs best in that constrained setting. We hypothesize that the added noise inhibits the backbone transformer from encoding fine-grained information into $\mathbf{z}^{s}_{\text{long}}$, limiting the MLP’s ability to reconstruct detailed audio. However, combining this with a short-context transformer computing $\mathbf{z}^{s}_{\text{short}}$ yields the best results (Tab. 5), likely because the clean short-term context restores local detail needed to model the distribution of the next $\mathbf{x}^{s}$.
At inference time, we omit noise augmentation and achieve our best results with this setup. Providing historical context to the diffusion head was used in the TTS model DiTAR (Jia et al., 2025), which reported a significant reduction in word error rate when using it.
4.4 Head Batch Multiplier
Training is bottlenecked by the cost of generating the conditioning variable $\mathbf{z}^{s}_{\text{long}}$ via the large causal transformer. To address this, we introduce the Head Batch Multiplier, which amortizes this cost by reusing $\mathbf{z}^{s}_{\text{long}}$ multiple times per training step. Specifically, for each input sequence, we compute $\mathbf{z}^{s}_{\text{long}}$ once and use it across $N$ loss computations, each with independently sampled noise levels $t$ and $\epsilon$. This not only improves efficiency but also stabilizes training by averaging the loss over multiple samples. As shown in Tab. 5 and Fig. 2, this leads to faster convergence and better final performance at comparable training cost.
4.5 Gaussian temperature sampling
Sampling strategies, such as temperature sampling, have a significant impact on generation quality in the discrete setup, particularly for speech. To replicate this behavior in the continuous domain, we introduce a sampling heuristic that results in similar gains.
Let $\bm{\epsilon}\sim\mathcal{N}(0,I)$ be a standard multivariate Gaussian random variable and let $\mathbf{x}=f(\bm{\epsilon})$, where $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a differentiable function. By the change of variables formula, the probability density of $\mathbf{x}$ is given by: $p(\mathbf{x})=p(\bm{\epsilon})\cdot\left|\det\left(\frac{\partial f(\bm{\epsilon})}{\partial\bm{\epsilon}}\right)\right|^{-1}$, where $\frac{\partial f(\bm{\epsilon})}{\partial\bm{\epsilon}}$ is the Jacobian matrix of $f$ at $\bm{\epsilon}$. In theory, to apply temperature sampling with temperature $\tau$ we should raise both terms to the power $1/\tau$, but the Jacobian term is untractable. In practice, we noticed that we can apply temperature only on the Gaussian noise sampling (which is equivalent to sampling from a Gaussian with variance $1/\sqrt{\tau}$) and get significant improvements in metrics.
The effects of gaussian temperature are further discussed in Section 5.3.
Table 1: Speech compression models. Our VAE is on par with the VQ-VAE on acoustic quality (MOSNeT) and outperforms it on semantic discriminability (ABX).
Model Type Dims / RVQ Frame Rate (Hz) Bitrate (kbit/s) MOSNet ($\uparrow$) ABX ($\downarrow$)
VQ-VAE (Mimi) 8 RVQ 12.5 1.1kbps 3.11 9.4%
VAE 32 dims 12.5 - 3.15 8.1%
5 Experiments and results
We now present experimental results on both speech and music continuation. We generate 27 seconds following a prompt of 3 seconds. All detailed hyperparameters are in the Appendix Sec. C. All metrics reported with confidence intervals use 95% confidence. Metrics without confidence intervals are metrics computed over the entire test dataset.
5.1 Speech Continuation
VAE: Our VAE is based on Mimi (Défossez et al., 2024b) but enforces gaussian inner latents instead of a categorical distribution. Like in Mimi, to enforce semanticity of the representations, we distill WavLM into the inner latent representation with a cosine similarity loss. Unlike Mimi, which applies this loss only to the first codebook, we extend it to the entire latent representation.
Model and dataset: Starting from Helium-1 (Kyutai, 2025), a pretrained 2B parameters multilingual text LM as backbone, we train on French and English speech data following Défossez et al. (2024b) to learn continuation. To enhance the stability and coherence of speech continuation, we adopt the concept of inner monologue (Défossez et al., 2024b)—a latent textual representation of the model’s own speech, aligned such that each word is positioned at the timestep corresponding to its spoken occurrence. This implies that, at each timestep $s$, the backbone transformer takes both text tokens and speech latents as input, and that its output $\mathbf{z}^{s}_{\text{long}}$ is passed through a linear layer which produces text logits alongside conditioning the consistency head. This internal text stream acts as a semantic scaffold, as it represents the next word to be pronounced, guiding the generation of audio tokens by grounding them in a linguistic form. Crucially, like in Défossez et al. (2024b), we introduce a temporal delay of 2 time steps (160ms) between the inner monologue and the corresponding audio tokens. This delay allows the model to access textual content prior to generating acoustic latents, decoupling high-level planning from low-level synthesis. For speech generation, we didn’t notice any gains from introducing a short context transformer and noising the latents before feeding them to the backbone, resulting in a simpler model architecture.
Results: Tab. 1 shows that our 32-dimensional VAE matches an 8-RVQ Mimi codec on MOSNet (Lo et al., 2019), which measures audio quality, and exceeds it on the ABX metric (Schatz et al., 2013). ABX evaluates phonetic discriminability by testing whether a word like “bat” is represented closer to another “bat” utterance than to a similar-sounding word like “bit”, based on latent distances.
Tab. 2 shows that the 1-step Consistency model outperforms the RQ-Transformer with 8 RVQ on all our automatic and human based metrics as well as on speed. For automatic metrics, we compute PPX and VERT Lakhotia et al. (2021). The PPX metric measures the semantic meaningfulness of the generated speech. To do so we generate 1000 excerpts of speech of 30 second, we use Whisper Radford et al. (2022) to compute textual transcriptions and finally compute the negative log-likelihood of the text tokens with a Mistral 7B LLM Jiang et al. (2023) and convert it to Perplexity. Because a model that generates poorly diverse but good quality sentences would perform well on the PPX metric, the authors of Lakhotia et al. (2021) introduce the VERT metric (for diVERsiTy) which is a geometric mean of self- and auto-BLEU metrics. We use the official implementation from the fairseq Ott et al. (2019) repository.
To assess perceptual quality, we conduct two human evaluation studies involving 50 participants and 50 randomly selected examples from the English test set. Each participant rates 10 examples across the following evaluation protocols: For Acoustic Quality, participants are presented with all model continuations for the same prompt, including the ground truth reference, and rate the acoustic quality of each continuation on a 1 to 5 scale. For Meaningfulness, participants are shown two continuations of the same prompt and select the one that is the most meaningful. These pairwise preferences are used to compute an Elo score (see Sec. B).
Notably, we note a clear quality and meaningfulness improvement with our temperature method. Given that there is a text stream to guide the audio generation, we expected the CALM to match the baseline on meaningfulness rather than outperforming it. This phenomenon could be due to less model capacity in the backbone being allocated to audio manipulation, allowing more for text prediction. On the inference speed side, the consistency head is $\times 12.3$ faster than the RQ-Transformer. The overall gain to perform a full inference of 30 seconds is $\times 1.3$.
Table 2: Comparison of speech continuation models: 8-RVQ RQ-transformer vs 1-step Consistency model head, with 2 temperature options.
Model Type Sampling Overall Sampler % Time in PPX ($\downarrow$) VERT ($\downarrow$) Acoustic Meaningfulness
temperature Speedup ($\uparrow$) Speedup ($\uparrow$) Sampler ($\downarrow$) Quality($\uparrow$) Elo ($\uparrow$) Rank ($\downarrow$)
Reference $20.2\pm 0.1$ $25.2\pm 0.1$ $4.02\pm 0.11$ $2180\pm 30$
RQ-transformer 8 RVQ 1.0 $\times 1.0$ $\times 1.0$ 26.7% $52.4\pm 0.2$ $36.3\pm 0.2$ $2.42\pm 0.12$ $1841\pm 25$ 4
RQ-transformer 8 RVQ 0.8 $\times 1.0$ $\times 1.0$ 26.7% $26.8\pm 0.1$ $33.1\pm 0.1$ $2.75\pm 0.14$ $1870\pm 30$ 3
CALM - Consistency - 1 step 1.0 $\times 1.3$ $\times 12.3$ 2.9% $42.9\pm 0.2$ $34.3\pm 0.2$ $2.82\pm 0.13$ $1947\pm 28$ 2
CALM - Consistency - 1 step 0.8 $\times 1.3$ $\times 12.3$ 2.9% $\bm{23.8}\pm\bm{0.1}$ $\bm{31.2}\pm\bm{0.1}$ $\bm{3.45}\pm\bm{0.14}$ $\bm{2023}\pm\bm{27}$ $\bm{1}$
5.2 Music Continuation
Dataset: We use a randomly selected subset of 400K songs (approximately 20K hours with 32kHz mono format) from the LAION-Disco-12M dataset, ensuring broad coverage across musical genres.
VAE: Our variational autoencoder (VAE) and codec architecture are adapted from the Mimi codec (Défossez et al., 2024b), originally designed for 24kHz speech at 12.5Hz. We trained it to compress 32kHz mono music with a 25Hz frame rate. We experiment with bottleneck sizes of 96 and 128 dimensions. For comparison, MusicGen’s EnCodec model (Copet et al., 2023) also operates at 32kHz but uses a 4-level RVQ at 50Hz. In Tab. 3, we report reconstruction metrics (audio ViSQOL Chinen et al. (2020) and SISNR), showing that a 32-dim VAE matches MusicGen’s codec, and that at least 96 dimensions are needed for our VAE to match the quality of the 32-level RVQ configuration.
Table 3: Music compression models. At least 96 VAE latent dimensions are required to outperform the 32-RVQ codec on reconstruction metrics. EnCodec has been retrained on our dataset.
Model Type Dims / RVQ Frame Rate (Hz) Bitrate (kbps) VisQOL ($\uparrow$) SISNR ($\uparrow$)
EnCodec Copet et al. (2023) 4 RVQ 50 2.2 2.41 5.62
VQ-VAE (inspired from Mimi) 32 RVQ 25 8.8 3.63 9.61
VAE 32 dims 25 2.23 5.51
VAE 96 dims 25 3.65 9.76
VAE 128 dims 25 4.01 10.3
Table 4: Comparison of music generation models across speed and quality metrics for 30 seconds generation. Consistency-based models provide up to a $2.2\times$ overall speedup and a $19.3\times$ sampler head speedup compared to the RQ-Transformer 32 RVQ baseline, while achieving improved FAD scores and equivalent human ratings. TrigFlow achieves the best qualitative results but has significantly higher inference time. Since MusicGen only uses a linear layer to sample its token we consider its inference cost as negligible.
Model Overall Sampler % time in FAD ($\downarrow$) Acoustic Enjoyment
Speedup ($\uparrow$) Speedup ($\uparrow$) Sampler ($\downarrow$) Quality ($\uparrow$) Elo ($\uparrow$) Rank ($\downarrow$)
Reference 3.84 $\pm$ 0.08 2166 $\pm$ 33 -
RQ-transformer 32 RVQ (baseline) $\times$ 1.0 $\times$ 1.0 57.7% 1.06 $\pm$ 0.06 2.85 $\pm$ 0.07 1824 $\pm$ 29 4
RQ-transformer 16 RVQ $\times$ 1.5 $\times$ 2.2 38.0% 1.43 $\pm$ 0.07 2.76 $\pm$ 0.07 1781 $\pm$ 29 5
CALM - Consistency - 1 Step $\times$ 2.2 $\times$ 19.3 6.6% 0.83 $\pm$ 0.04 2.90 $\pm$ 0.07 1857 $\pm$ 28 2
CALM - Consistency - 4 Steps $\times$ 1.9 $\times$ 5.4 20.1% 0.71 $\pm$ 0.05 3.07 $\pm$ 0.07 1847 $\pm$ 24 3
CALM - TrigFlow - 100 Steps $\times$ 0.3 $\times$ 0.2 86.6% 0.64 $\pm$ 0.04 3.12 $\pm$ 0.07 1921 $\pm$ 29 1
MusicGen Medium $\times$ 1.3 0.0% 1.72 $\pm$ 0.12 2.62 $\pm$ 0.07 1761 $\pm$ 33 6
Table 5: Ablation study on music CALM Consistency 4-steps model components, after 250K training steps. Removing noise augmentation or the short-context transformer leads to significant performance drops. Final row approximates the MAR configuration from Li et al. (2024).
Model Variant FAD ($\downarrow$)
Base (CALM - Consistency - 4 steps) 0.93 $\pm$ 0.06
   w/o Head Batch Multiplier 1.32 $\pm$ 0.09
   w/o Noise Augmentation 1.63 $\pm$ 0.11
   w/o Short Context Transformer 4.03 $\pm$ 0.16
   w/o Any of the above 8.38 $\pm$ 0.17
Model: Our music CALM builds on the MusicGen Medium backbone, a 1.35B parameter Transformer (see Sec. C for all the hyperparameters). We compare our method against: (1) Two discrete models using 32 and 16 RVQ codecs, each employing the same 1.35B backbone and a RQ-transformer for parallel codes prediction; (2) a MusicGen Medium variant trained with EnCodec without textual conditioning, using the same backbone and a delay pattern for codebook interleaving. Both MusicGen Medium and EnCodec models were trained on our dataset.
Results and Ablation: In Tab. 4, we report both objective metrics and the results of a human evaluation study for the task of music continuation, conditioned on a 3-second prompt. We compute the speed-up compared to the RQ-Transformer 32 RVQ, the Fréchet Audio Distance (FAD) which is the distance between Gaussian distributions fitted on VGG-obtained embeddings of the real and generated samples. We compute it on 4,000 model-generated continuations from the test set. The Acoustic Quality is a MOS score between 1 and 5. The Enjoyment metric is an Elo score (see Sec. B), computed by making human rater choose their favorite music out of generated pairs with the same 3s prompt. The evaluation protocol uses the same amount of annotators and examples as the speech study. We observe that CALM with consistency outperforms the 32 RVQ RQ-Transformer baseline on computed and human metrics while being $\times 1.9$ to $\times 2.2$ times faster for the overall speedup. While the RQ-Transformer takes $57.7\%$ of the inference time for the baseline, the consistency head only takes $6.6\%$ to $20.1\%$. As well, we train a CALM model with a TrigFlow head instead of consistency and it outperforms all the models but to the price of a slow inference.
Figure 2: Effect of the head batch multiplier value. Training a model (music consistency CALM) with a higher batch size multiplier fastens the convergence for the FAD metric. All the evaluations are done with 4 steps of consistency at inference time.
An ablation study (Tab. 5) on music CALM Consistency 4 steps shows the importance of each component. The experiments are run for 250K steps, which explains that the base model’s FAD is worse than the one reported in Tab. 4. The final row that is the closest to the MAR framework (consistency replacing diffusion) fails to produce high-quality music. In Fig. 2, we show that the FAD decreases much faster over time when we train consistency CALM models with a bigger head batch multiplier. All evaluations are done with 4 steps of consistency. We keep the value of 8 for all of our experiments as a higher value would lead to out of memory issues at training time.
Scalability: We show in Sec. A.2 that CALM does improve with a bigger 3B parameters backbone. However, we leave a complete scalability study for future work.
Necessity of Consistency for a fast inference: We show in Tab. 6 that the consistency framework largely outperforms the TrigFlow framework for 10 inference steps and less (the regime where the Real time factor is greater than 1).
Table 6: Generation efficiency of TrigFlow CALM and Consistency CALM models. We compute the inference time to generate 30 seconds of audio, the corresponding Real Time Factor (RTF) as well as the FAD metric for different numbers of inference steps. For less than 10 steps (i.e. when streaming applications are possible, the consistency method outperforms the TrigFlow method).
# Steps Time (s) RTF FAD (TrigFlow) FAD (Consistency)
1 16.7 $\times 1.80$ $0.83\pm 0.04$
4 20.4 $\times 1.47$ $28.83\pm 0.20$ $0.71\pm 0.05$
10 27.7 $\times 1.08$ $4.62\pm 0.07$ $0.73\pm 0.05$
25 46.1 $\times 0.65$ $0.79\pm 0.04$ $0.96\pm 0.06$
50 76.8 $\times 0.39$ $0.74\pm 0.05$ $1.46\pm 0.06$
100 136.4 $\times 0.22$ $0.64\pm 0.04$ $2.05\pm 0.07$
5.3 Effect of gaussian temperature sampling
We evaluate how our proposed gaussian temperature sampling method affects acoustic diversity for consistency models, in comparison with temperature sampling on its discrete counterpart. More precisely, we compute WavLM speaker embeddings (Chen et al., 2021) over 100 unprompted speech generations with different values of temperature, both for an RQ-transformer and CALM. In both cases, the inner monologue text stream is generated with 0.8 temperature to ensure text diversity. We then average the pairwise cosine similarities of these embeddings, as a measure of diversity: the higher the average similarity, the lower speaker diversity in generated audio. The reference speaker similarity number is computed over 100 examples from the ground truth dataset. Fig. 3 shows, as expected, that speaker similarity tends to decrease with temperature, which means that temperature increases diversity with a similar trajectory to the discrete models.
Figure 3: Average pairwise speaker similarity over 100 unprompted 10s generations, or 100 10s examples from the ground truth dataset as reference. As expected, for both methods models generate more diverse speakers (i.e. less pairwise speaker similarity) as temperature increases.
6 Conclusion
We present Continuous Audio Language Models (CALM), a novel framework for autoregressive audio generation that operates directly in the continuous latent space of a VAE, bypassing the limitations of discrete quantization. Replacing RVQ or diffusion heads with consistency models significantly reduces inference cost while improving sample quality as shown by our experiments. Our architecture combines noise-injected long-term context and clean short-term context, implemented via a dual-transformer design. We introduce practical innovations such as temperature sampling and a head batch multiplier to further improve sampling quality and training efficiency. We demonstrate the effectiveness of our approach across both speech and music generation tasks. Our results suggest that continuous modeling offers a compelling alternative to discrete tokenization for high-quality, efficient, and scalable autoregressive audio generation.
Appendix A Supplementary Experiments
A.1 Short-context transformer window
To study the influence of the context window $K$ of the short-context transformer, we performed an hyperparameter search with $K=5,10,20,40$ and trained our models for 500K training steps. We observe that finding an optimal value of $K$ is not critical even though it seems that 5 is not enough.
Table 7: FAD after 500K training steps for different short-context Transformer contexts $K$.
$K$ in short-context Transformer FAD
5 $0.85\pm 0.05$
10 $0.76\pm 0.04$
20 $0.73\pm 0.04$
40 $0.78\pm 0.04$
A.2 Scalability of CALM
In order to see if all the hyperparameters of our CALM method transfer well to more model parameters, we trained a consistency music CALM model with a larger 3B backbone (with a model dimension of 2048, 32 heads and 48 layers) as well as a discrete RQ-Transformer with 32-RVQ model with the same backbone size. For the CALM model, we use 4 consistency steps at inference. The results are as follow (we compare with the main model that has a 1.3B backbone):
Table 8: Scalability of the backbone for CALM and RQ-Transformer methods.
Model FAD
CALM with 3B backbone $0.62\pm 0.05$
32 RVQ RQ-Transformer with 3B backbone $0.98\pm 0.06$
CALM with 1.3B backbone $0.71\pm 0.05$
32 RVQ RQ-Transformer with 1.3B backbone $1.06\pm 0.06$
Appendix B Human evaluation methods
Audio clips are always 30s second total, with a 3s prompt coming from a ground truth audio. Each experiments has 50 samples for each method. There are 50 raters. Each of them sees 10 samples. Raters were payed £9 / hour for their contribution.
B.1 Speech continuation
Acoustic quality assessment: How would you rate the overall quality of this audio clip? Consider aspects such as clarity, balance, richness, and naturalness. Listen to at least 10 seconds of audio before deciding. 1 clip is presented, possibilities are bad, poor, fair, good, excellent.
Meaningfulness: Which of these two audio clips feels more like meaningful and natural speech? The first 3 seconds are identical. Listen to at least 10 seconds of each clip. 2 clips are presented, ties are possible. Elo scores are Bayesian estimates of the posterior mean in a Bradley-Terry model.
B.2 Music continuation
Acoustic quality assessment: How would you rate the overall quality of this music? Consider aspects such as clarity, balance, richness, and naturalness. Listen to at least 10 seconds of audio before deciding. 1 clip is presented, possibilities are bad, poor, fair, good, excellent.
Music enjoyment: Which music do you enjoy listening to more? The first 3 seconds are identical. Listen to at least 10 seconds of each clip. 2 clips are presented, ties are possible. Elo scores are Bayesian estimates of the posterior mean in a Bradley-Terry model.
B.3 Bayesian Elo Score
The Meaningfulness metric for speech and the Enjoyment metric for music are both Bayesian Elo Scores. Elo score is used to rank models based on some pairwise comparisons of audio samples. Given two models $A$ and $B$, the probability that $A$ is preferred over $B$ is:
$P(A>B)=\frac{1}{1+10^{(E_{B}-E_{A})/400}}$ (4)
where $E_{A}$ and $E_{B}$ are the Elo scores of each model. Unlike a traditional Elo score, the Bayesian Elo score uses a Gamma prior, so that one can derive confidence intervals over the posterior distribution.
By defining $S_{A}$ such as $E_{A}=400\log_{10}(S_{A})+c$ with $c$ being a constant, we obtain:
$P(A>B)=\frac{S_{A}}{S_{A}+S_{B}}$ (5)
which is a Bradley-Terry (Bradley & Terry, 1952) model. There are a few different methods to estimate the parameters of a Bradley-Terry model. We use the iterative one from (Caron & Doucet, 2010) where $S_{A}^{0}$ follows a Gamma prior with parameters $\alpha^{0},\beta^{0}$. By denoting $w_{A}$ the number of times where method $A$ won against any other methods and $n_{AB}$ the number of times where $A$ and $B$ are compared, $S_{A}^{t}$ is computed with the following update rule until convergence:
$S_{A}^{t+1}=\frac{\alpha+w_{A}}{\beta+\sum_{B\neq A}\frac{n_{AB}}{S_{A}^{t}+S_{B}^{t}}}$ (6)
which is the mean of the Gamma distribution with updated parameters $\alpha_{A}^{t+1},\beta_{A}^{t+1}$:
$\alpha_{A}^{t+1}=\alpha^{0}+w_{A},\qquad\text{and}\qquad\beta_{A}^{t+1}=\beta^{0}+\sum_{B\neq A}\frac{w_{A,B}}{S_{A}^{t}+S_{B}^{t}}.$ (7)
Iterating over $t$ allows to reach a fix point, we run 30 of them, once we have collected all the pairs. We use $\alpha=0.1,\beta=0.1,c=2000$ so that in absence of any data, $S_{A}=1$ and $E_{A}=2000$. Confidence interval are 95% confidence interval according to the posterior (the 2.5th and 97.5th percentiles).
Appendix C Hyperparameters
We present in Tab. 10 the parameters of the music and speech VAE used for CALMs. For CALMs and the RQ-Transformer based discrete LMs, the hyperparameters are presented in Tab. 9.
Table 9: Model and training hyperparameters
Music continuation Speech continuation
Backbone Transformer
Model dimension 1536 2560
MLP dimension 6336 10560
Number of heads 24 20
Number of layers 48 24
Learning rate $1\cdot 10^{-4}$ $5\cdot 10^{-5}$
Number of parameters 1.35B 2.2B
RQ-Transformer head (RVQ model)
Model dimension 1024 1024
MLP dimension 4096 4096
Number of heads 16 16
Number of layers 6 6
Number of parameters 701M 701M
Consistency sampler head (ours)
Number of layers 12 6
MLP dimension 3072 512
Gating SiLU SiLU
Number of parameters 601M 10M
Short context transformer (ours)
Model dimension 1536
MLP dimension 6336
Number of heads 24
Number of layers 4
Context 10
Number of parameters 113M
Audio embedding manipulations
Center and normalize
Noise before entering backbone
Training parameters
Head batch multiplier 8 8
Optimizer AdamW $\beta_{1}=0.9,\beta_{2}=0.95$ AdamW $\beta_{1}=0.9,\beta_{2}=0.95$
Batch size 48 144
Audio sample length 30s 300s
LR Schedule cosine cosine
Learning rate $1\cdot 10^{-4}$ $2\cdot 10^{-4}$
GPU used 16 H100 48 H100
Number of training steps 500k 150k
Initial checkpoint Helium-1 (Kyutai, 2025)
Inner monologue
Acoustic delay - 2 frames (160 ms)
Table 10: VAE hyperparameters
Music VAE Speech VAE
General
Sample rate 32kHz 24kHz
Frame rate 25Hz 12.5Hz
Latent dimension 128 32
Architecture
Convolutions ratios 8, 8, 5, 4 6, 5, 4, 4, 4
Num transformer encoder layers 4 8
Num transformer decoder layers 4 8
Transformer context 30s 10s
Training parameters
Batch size 64 64
Audio sample length 12s 12s
KL loss weight 0.01 0.01
Reconstruction loss
Distillation loss weight 25
LR Schedule cosine cosine
Learning rate $8\cdot 10^{-4}$ $8\cdot 10^{-4}$

### v3

@@
 Experiments on speech and music demonstrate improved efficiency and fidelity over state-of-the-art discrete audio language models, facilitating lightweight, high-quality audio generation.
-Samples are available at hf.co/spaces/kyutai/calm-samples.
+Samples are available at iclr-continuous-audio-language-models.github.io.
+Finally, we release Pocket TTS, an open-source 100M-parameter text-to-speech model that can run faster than real time on a laptop CPU: github.com/kyutai-labs/pocket-tts.
 1 Introduction
@@
 Standard variational autoencoders (VAEs) are easier to train, are not affected by issues such as codebook collapse, and can reconstruct audio at higher fidelity for the same latent dimensionality.
-Pioneering work in the vision domain that autoregressively models continuous sequences includes GIVT (Tschannen et al., 2024) and MAR (Li et al., 2024), followed by some larger models in the image domain (Fan et al., 2025; Gu et al., 2025) and attempts in the audio domain (Turetzky et al., 2024; Pasini et al., 2024b; Jia et al., 2025).
+Pioneering work in the vision domain that autoregressively models continuous sequences includes GIVT (Tschannen et al., 2024) and MAR (Li et al., 2024), followed by some larger models in the image domain (Fan et al., 2025; Gu et al., 2025) and attempts in the audio domain (Turetzky et al., 2024; Pasini et al., 2024b; Jia et al., 2025; Yang et al., 2025).
 In MAR, the authors model the per-token probability distribution with a diffusion model (in the form of a small MLP) conditioned on a latent variable modeled by an autoregressive transformer backbone.
 SALAD (Turetzky et al., 2024) and DiTAR (Jia et al., 2025) adapt MAR-style diffusion heads for text-to-speech modeling, achieving better audio quality than discrete baselines.
-However, these works are limited to text-to-speech on small-scale and domain-specific datasets, leaving open the question of how well they can adapt to richer audio domains such as music.
-In this work, we propose Continuous Audio Language Models (CALM) that predict sequences in the latent space of a VAE, bypassing the need for quantization.
-While we build on the MAR architecture where a transformer backbone uses $(\mathbf{x}^{1},\ldots,\mathbf{x}^{s-1})$ to predict an intermediate latent $\mathbf{z}^{s}$, which then conditions a head (diffusion model) that models $p(\mathbf{x}^{s}|\mathbf{z}^{s})$, we find that without further improvements, it fails to generate rich audio content.
+However, these works are limited to text-to-speech on small-scale and domain-specific datasets, leaving open the question of how well they can adapt to more complex tasks such as speech continuation (without text supervision) and richer audio domains such as music.
+We apply CALM to 4 tasks which are speech and music continuation as well as text-to-speech and text-to-music.
+We propose Continuous Audio Language Models (CALM) that predict sequences in the latent space of a VAE, bypassing the need for quantization.
+While we build on the MAR architecture where a transformer backbone uses $(\mathbf{x}^{1},\ldots,\mathbf{x}^{s-1})$ to predict an intermediate latent $\mathbf{z}^{s}$, which then conditions a head (diffusion model) that models $p(\mathbf{x}^{s}|\mathbf{z}^{s})$, we find that without further improvements, it fails to generate rich audio content and is slow to sample from.
 To overcome this, we introduce several key innovations:
@@
 2.
-Diffusion-to-Consistency replacement: We replace the diffusion model with a continuous consistency model (Lu & Song, 2025) during training, significantly accelerating inference without compromising sample quality.
+Diffusion-to-Consistency replacement: We replace the diffusion model with a continuous consistency model (Lu and Song, 2025) during training, significantly accelerating inference without compromising sample quality.
 This change reduces the inference time of the sampler head by a factor of up to $\times 20$ in our music experiments and $\times 12$ in our speech experiments compared to the use of a model that uses an RQ-Transformer as head.
-This method is modality-agnostic and may benefit other domains such as image generation.
 3.
@@
 Head batch multiplier: Sampling multiple noise levels at training time for the same latent highly accelerates training for a small cost.
-Our model generates high-quality speech and music with lower inference time, outperforming state-of-the-art discrete-token autoregressive models in both efficiency and audio fidelity.
-To our knowledge, this is the first application of autoregressive continuous latent modeling to both domains.
+5.
+Latent Classifier Free Guidance: we apply Classifier Free Guidance to the latent variable conditioning the consistency head for the conditioned CALM.
+6.
+Latent Distillation: once that we have chosen a latent CFG coefficient, we can distill the CFG computation of the backbone into a student backbone while keeping the same sampling head (MLP), hence dividing the batch size by 2 at inference time.
+As well, we can apply the distillation to a much smaller student backbone.
+Finally, by using these innovations, we introduce Pocket TTS which is a 100M-parameters text-to-speech model that can run faster than real-time on a laptop CPU.
+We detail the results of Pocket TTS in Sec.F and in the following technical report: kyutai.org/pocket-tts-technical-report.
 Figure 1: Overview of our model.
@@
 Autoregressive audio language models.
-Early autoregressive audio models operated on raw waveforms such as WaveNet (van den Oord et al., 2016) or learned discrete codes via VQ-VAE, such as in Jukebox (Dhariwal et al., 2020).
+Early autoregressive audio models operated on raw waveforms such as WaveNet (van den Oord et al., 2016), learned discrete codes via VQ-VAE, such as in Jukebox (Dhariwal et al., 2020) or continuous word-sized audio tokens (Algayres et al., 2023).
 The advent of neural audio codecs (Zeghidour et al., 2021; Défossez et al., 2024a) enabled high-quality vector-quantized tokenizers for general audio.
@@
 (2024) use a transformer to autoregressively model the latent space of a VAE trained on images, parameterize a Gaussian Mixture Model, and train it with a cross-entropy loss.
-The authors of MAR (Li et al., 2024) obtain better results by replacing the Gaussian mixture model with a diffusion-model enabling the approximation of more flexible distributions.
+The authors of MAR (Li et al., 2024) obtain better results by replacing the Gaussian mixture model with a diffusion-model enabling the approximation of more diverse distributions.
 In MAR, a large transformer backbone predicts a continuous embedding $\mathbf{z}^{s}$ given $(\mathbf{x}^{1},\ldots,\mathbf{x}^{s-1})$, which then conditions a small MLP diffusion network that models the probability distribution $p(\mathbf{x}^{s}|\mathbf{z}^{s})$ of the next latent.
@@
 They reduce the number of diffusion steps from 100 to 8 with a similar image quality.
-Apart from this work, we are not aware of other works that try to reduce the inference time of the diffusion process in continuous autoregressive models.
 Remarkably, CALM achieves a quality comparable to the best discrete models with only one step of consistency modeling.
@@
 The authors observe that providing local context through patching was determinant to improve their model, which corroborates what we observe with the introduction of a short context transformer into our model.
-While these two studies focus on TTS modeling, we extend the method to music.
 In Pasini et al.
@@
 Finally, Music2Latent2 (Pasini et al., 2025) explores combining autoregressive and consistency modeling but for compression.
+IMPACT (Huang et al., 2025) explores MAR decoding for text-to-audio generation and sets a new standard for short latency models on the AudioCaps (Kim et al., 2019) benchmark.
+More recently, MingUni-Audio (Yan et al., 2025) shows that continuous speech language models can scale to 20B Mixture of Experts models with 3B active parameters.
+Some other autoregressive speech models work in the continuous domain thanks to spectral representations such as MELLE (Meng et al., 2024) for TTS and Flow-Omni (Yuan et al., 2024) for speech-to-speech conversation.
 3 Background
@@
 This approach enables efficient parallel modeling of RVQ sequences, allowing all codebooks corresponding to a single timestep to be generated simultaneously without introducing any delay.
-However, a key limitation lies in the use of the RQ-Transformer, which is computationally intensive; as its resource requirements scale with the number of RVQ.
+However, a key limitation lies in the use of the RQ-Transformer, which is computationally intensive; as its resource requirements scale with the number of RVQ, or even with the square of the number of RVQ, if the attention dominates the computation cost.
 3.2 Consistency Models
@@
 To ensure correct behavior, $f_{\phi}$ must satisfy the boundary condition $f_{\phi}(\mathbf{x},0)=\mathbf{x}$, thus leading to the common parameterization $f_{\phi}(\mathbf{x}_{t},t)=c_{\text{skip}}(t)\mathbf{x}_{t}+c_{\text{out}}(t)F_{\phi}(\mathbf{x}_{t},t)$, where $F_{\phi}$ is a neural network and the coefficients satisfy $c_{\text{skip}}(0)=1$ and $c_{\text{out}}(0)=0$ to fulfill the boundary condition.
-By using $T=\frac{\pi}{2}$ and $\alpha_{t}=\cos(t),\sigma_{t}=\sin(t)$, Lu & Song (2025) derive the following continuous-time consistency loss where $w_{\psi}(t)$ is an adaptive weighting function:
+By using $T=\frac{\pi}{2}$ and $\alpha_{t}=\cos(t),\sigma_{t}=\sin(t)$, Lu and Song (2025) derive the following continuous-time consistency loss where $w_{\psi}(t)$ is an adaptive weighting function:
 $\mathcal{L}_{\text{CM}}(\phi,\psi)=\mathbb{E}_{\mathbf{x}_{t},t}\left[\!\frac{e^{w_{\psi}(t)}}{D}\left\|F_{\phi}\left(\mathbf{x}_{t},t\right)-F_{\phi^{-}}\left(\mathbf{x}_{t},t\right)-\cos(t)\frac{df_{\phi^{-}}(\mathbf{x}_{t},t)}{dt}\right\|_{2}^{2}-w_{\psi}(t)\right].$ (1)
+Lagrangian Self-Distillation.
+We also explore a new 1-step flow-matching method named Latent Self-Distillation (LSD) that has been introduced in Boffi et al.
+(2025).
+In this paper, the authors unify most 1-step methods into 2 categories and develops a third one (LSD) which appears to be more stable at training.
+See Sec.
+A for the equations.
 3.3 Autoregressive Modeling of Continuous Latents via Diffusion.
@@
 While training the model with adversarial losses and VAE regularization without any reconstruction losses improves the quality of the model for speech, it degrades the reconstruction quality for music.
-Semantic distillation is performed for the speech VAE similarly as in Mimi, with WavLM (Chen et al., 2021) as teacher.
+Semantic distillation is performed for the speech VAE similarly as in Mimi, with WavLM (Chen et al., 2021b) as teacher.
 There is no semantic distillation for the music model and we let this for future work as semantic content is harder to define for music.
 The loss is:
-$\mathcal{L}_{\text{VAE}}=\lambda_{\text{t}}\mathcal{L}_{\text{t}}(x,\hat{x})+\lambda_{\text{f}}\mathcal{L}_{\text{f}}(x,\hat{x})+\lambda_{\text{adv}}\mathcal{L}_{\text{adv}}(\hat{x})+\lambda_{\text{feat}}\mathcal{L}_{\text{feat}}(x,\hat{x})+\lambda_{\text{KL}}\mathcal{L}_{\text{KL}}$ (2)
-where $\mathcal{L}_{\text{t}}$ and $\mathcal{L}_{\text{f}}$ are the temporal and frequential reconstruction losses, $\mathcal{L}_{\text{adv}}$ is the adversarial loss, $\mathcal{L}_{\text{feat}}$ is the feature matching loss and $\mathcal{L}_{\text{KL}}$ is the KL regularization applied to the VAE bottleneck.
-4.2 Consistency-Based Continuous Audio Language Model (CALM)
+$\mathcal{L}_{\text{VAE}}=\lambda_{\text{t}}\mathcal{L}_{\text{t}}(x,\hat{x})+\lambda_{\text{f}}\mathcal{L}_{\text{f}}(x,\hat{x})+\lambda_{\text{adv}}\mathcal{L}_{\text{adv}}(\hat{x})+\lambda_{\text{feat}}\mathcal{L}_{\text{feat}}(x,\hat{x})+\lambda_{\text{KL}}\mathcal{L}_{\text{KL}}+\lambda_{\text{distill}}\mathcal{L}_{\text{distill}}$ (2)
+where $\mathcal{L}_{\text{t}}$ and $\mathcal{L}_{\text{f}}$ are the temporal and frequential reconstruction losses, $\mathcal{L}_{\text{adv}}$ is the adversarial loss, $\mathcal{L}_{\text{feat}}$ is the feature matching loss, $\mathcal{L}_{\text{KL}}$ is the KL regularization applied to the VAE bottleneck, and $\mathcal{L}_{\text{distill}}$ is the WavLM distillation loss applied for the speech VAE.
+4.2 Continuous Audio Language Model (CALM) architecture
 Let $\bigl(\mathbf{x}^{1},\ldots,\mathbf{x}^{S}\bigr)$ denote the sequence of continuous latent vectors produced by a VAE encoder.
 As illustrated in Fig.
-1, our model comprises three main components:
+1, our model comprises three main components.
+The motivations behind these design choices are described in Sec.
+4.3 and the ablation study in Tab.
+5 justifies these choices.
 1. Causal Backbone Transformer with Noise Injection
-We build on the MAR framework (Li et al., 2024) by employing a causal Transformer $T_{\text{long},\theta^{1}}$ to capture long-term dependencies.
-During training, each input latent $\mathbf{x}^{s}$ is replaced with a noised version $\tilde{\mathbf{x}}^{s}$ (see Sec.
-4.3 and Ablation 5).
+We build on the MAR framework (Li et al., 2024) by employing a causal Transformer $T_{\text{long},\theta^{1}}$ to capture long‑term dependencies.
+However, during preliminary experiments, we realized that music generation models with the MAR framework were generating poor quality audio and diverging quickly at inference because they were not robust to error accumulation.
+Pasini et al.
+(2024b) introduced a noise augmentation trick at training time in order to tackle this problem.
+Given a sequence $(\mathbf{x}^{1},\mathbf{x}^{2},\ldots,\mathbf{x}^{S})$, they sample $k_{s}\sim\mathcal{U}(0,1)$ and $\bm{\epsilon}_{s}\sim\mathcal{N}(0,\mathbf{I})$ for every $s\in\{1,\ldots,S\}$ and use a noised input to the backbone $\tilde{\mathbf{x}}^{s}=k_{s}\bm{\epsilon}_{s}+(1-k_{s})\mathbf{x}^{s}$ for every $s$.
+Early experiments showed that preserving the variance of $\tilde{\mathbf{x}}^{s}$ improved quality, so that we use instead $\tilde{\mathbf{x}}^{s}=\sqrt{k_{s}}\bm{\epsilon}_{s}+\sqrt{1-k_{s}}\mathbf{x}^{s}$.
 We don’t perform any noise augmentation at inference time.
-Concretely, $\mathbf{z}_{\text{long}}^{s}\;=\;T_{\text{long},\theta^{1}}\bigl{(}\tilde{\mathbf{x}}^{1},\,\ldots,\,\tilde{\mathbf{x}}^{s-1}\bigr{)}$.
+Thus, $\mathbf{z}_{\text{long}}^{s}\;=\;T_{\text{long},\theta^{1}}\bigl(\tilde{\mathbf{x}}^{1},\,\ldots,\,\tilde{\mathbf{x}}^{s-1}\bigr)$.
 Noise injection prevents error accumulation during inference, but as shown in Tab.
-5 is insufficient alone for high-quality music generation.
-2. Short-Context Transformer
-To supply local, high-resolution context to the denoising head, we introduce a lightweight causal Transformer that attends the $K$ previous clean latents (we use $K=10$, $\sim$ 0.4s of music): $\mathbf{z}_{\text{short}}^{s}\;=\;T_{\text{short},\theta^{2}}\bigl{(}\mathbf{x}^{s-K},\,\ldots,\,\mathbf{x}^{s-1}\bigr{)}$.
-This short-context embedding $\mathbf{z}_{\text{short}}^{s}$ supplies fine-grained information potentially lost through noise injection in the backbone.
-We show in Tab.
-7 that the value of K is not a decisive hyperparameter.
-3. Consistency-Model Head
-Finally, a small MLP-based consistency model $f_{\phi}$ is conditioned on the sum of the long-term and short-term features, $\mathbf{Z}^{s}=\mathbf{z}_{\text{long}}^{s}+\mathbf{z}_{\text{short}}^{s}$.
-At inference time, for 1-step, the next latent $\hat{\mathbf{x}}^{s}$ is sampled through: $\bm{\epsilon}\sim\mathcal{N}(0,I),t=1$, $\hat{\mathbf{x}}^{s}\;=\;f_{\phi}\bigl{(}\mathbf{x}^{s}_{1}=\bm{\epsilon},t=1,\mathbf{Z}^{s}\bigr{)}$.
-In addition to consistency, we experiment with the TrigFlow (Lu & Song, 2025) formulation of flow-matching for the MLP.
-Although TrigFlow yields marginally higher fidelity, its inference cost makes it impractical for real-time use.
-This tradeoff is studied in Sec.
-5.2, Tab.
-4.
-Together, these three components form a continuous autoregressive model that (i) leverages noise-robust long-term modeling, (ii) preserves local detail via short-context conditioning, and (iii) achieves rapid, high-fidelity latent sampling through consistency modeling.
+5 is insufficient alone for high‑quality music generation.
+2. Short‑Context Transformer
+To supply local, high‑resolution context to the denoising head, we introduce a lightweight causal Transformer that attends the $K$ previous clean latents (we use $K=10$, $\sim$ 0.4s of music): $\mathbf{z}_{\text{short}}^{s}\;=\;T_{\text{short},\theta^{2}}\bigl(\mathbf{x}^{s-K},\,\ldots,\,\mathbf{x}^{s-1}\bigr)$.
+This short‑context embedding $\mathbf{z}_{\text{short}}^{s}$ supplies fine‑grained information potentially lost through noise injection in the backbone.
+We show in Sec.
+E.2 that the value of K is not a decisive hyperparameter but Tab.
+5 indicates that the short-context transformer is crucial to good quality generation.
+3. Consistency‑Model Head
+Finally, a small MLP-based consistency model $f_{\phi}$ is conditioned on the sum of the long‑term and short‑term features, $\mathbf{Z}^{s}=\mathbf{z}_{\text{long}}^{s}+\mathbf{z}_{\text{short}}^{s}$.
+At inference time, for 1-step generation, the next latent $\hat{\mathbf{x}}^{s}$ is sampled through: $\bm{\epsilon}\sim\mathcal{N}(0,I),t=1$, $\hat{\mathbf{x}}^{s}\;=\;f_{\phi}\bigl(\mathbf{x}^{s}_{1}=\bm{\epsilon},t=1,\mathbf{Z}^{s}\bigr)$.
+In addition to consistency, we experiment with the TrigFlow (Lu and Song, 2025) formulation of flow‑matching for the MLP.
+Although TrigFlow yields marginally higher fidelity, its inference cost makes it impractical for real‑time use.
+While Tab.
+4 shows it, this tradeoff is studied in Sec.
+E.3.
+Together, these three components form a continuous autoregressive model that (i) leverages noise‑robust long‑term modeling, (ii) preserves local detail via short‑context conditioning, and (iii) achieves rapid, high‑fidelity latent sampling through consistency modeling.
 The training objective for one sequence $\bigl(\mathbf{x}^{1},\ldots,\mathbf{x}^{S}\bigr)$ is defined by:
-$\small\mathcal{L}_{\text{CALM}}(\theta,\phi,\psi)=\sum_{s=1}^{S}\mathbb{E}_{t,\bm{\epsilon}}\left[\!\frac{e^{w_{\psi}(t)}}{D}\left\|F_{\phi}\left(\mathbf{x}^{s}_{t},t,\mathbf{Z}^{s}\right)-F_{\bar{\phi}}\left(\mathbf{x}^{s}_{t},t,\mathbf{Z}^{s}\right)-\cos(t)\frac{df_{\bar{\phi}}(\mathbf{x}^{s}_{t},t,\mathbf{Z}^{s})}{dt}\right\|_{2}^{2}-w_{\psi}(t)\right],$ (3)
+$\mathcal{L}_{\text{CALM}}(\theta,\phi,\psi)=\sum_{s=1}^{S}\mathbb{E}_{t,\bm{\epsilon}}\left[\!\frac{e^{w_{\psi}(t)}}{D}\left\|F_{\phi}\left(\mathbf{x}^{s}_{t},t,\mathbf{Z}^{s}\right)-F_{\bar{\phi}}\left(\mathbf{x}^{s}_{t},t,\mathbf{Z}^{s}\right)-\cos(t)\frac{df_{\bar{\phi}}(\mathbf{x}^{s}_{t},t,\mathbf{Z}^{s})}{dt}\right\|_{2}^{2}-w_{\psi}(t)\right],$ (3)
 where $\mathbf{Z}^{s}=\mathbf{z}_{\text{long}}^{s}+\mathbf{z}_{\text{short}}^{s}=T_{\text{long},\theta^{1}}\bigl(\tilde{\mathbf{x}}^{1},\,\ldots,\,\tilde{\mathbf{x}}^{s-1}\bigr)+T_{\text{short},\theta^{2}}\bigl(\mathbf{x}^{s-K},\,\ldots,\,\mathbf{x}^{s-1}\bigr)$,
@@
 4.3 Combining noisy long-term context and clean short-term context
-In preliminary experiments, we realized that the music generation model was diverging quite quickly during inference because it was not robust to error accumulation.
-Pasini et al.
-(2024b) introduce a noise augmentation trick at training time in order to tackle this problem.
-Given a sequence $(\mathbf{x}^{1},\mathbf{x}^{2},\ldots,\mathbf{x}^{S})$, they sample $k_{s}\sim\mathcal{U}(0,1)$ and $\bm{\epsilon}_{s}\sim\mathcal{N}(0,\mathbf{I})$ for every $s\in\{1,\ldots,S\}$ and compute $\tilde{\mathbf{x}}^{s}=k_{s}\bm{\epsilon}_{s}+(1-k_{s})\mathbf{x}^{s}$ for every $s$.
-After comparison and to preserve the variance of the $\mathbf{x}^{s}$ that are gaussians, we modify the noising formula as follows: $\tilde{\mathbf{x}}^{s}=\sqrt{k_{s}}\bm{\epsilon}_{s}+\sqrt{1-k_{s}}\mathbf{x}^{s}$
-In our case, applying this noising method to music generation slightly improves model stability but often reduces detail and instrument diversity, typically preserving only rhythmic elements, with audio fading into silence after 10–15 seconds.
+In preliminary experiments, music generation models trained with the MAR framework were diverging quickly during inference because they were not robust to error accumulations.
+Applying noise injection during training slightly improves model stability but often reduces detail and instrument diversity, typically preserving only rhythmic elements, with audio fading into silence after 10–15 seconds.
 Since Pasini et al.
@@
 5), likely because the clean short-term context restores local detail needed to model the distribution of the next $\mathbf{x}^{s}$.
-At inference time, we omit noise augmentation and achieve our best results with this setup.
-Providing historical context to the diffusion head was used in the TTS model DiTAR (Jia et al., 2025), which reported a significant reduction in word error rate when using it.
 4.4 Head Batch Multiplier
@@
 5 and Fig.
-2, this leads to faster convergence and better final performance at comparable training cost.
+3, this leads to faster convergence and better final performance at comparable training cost.
 4.5 Gaussian temperature sampling
 Sampling strategies, such as temperature sampling, have a significant impact on generation quality in the discrete setup, particularly for speech.
-To replicate this behavior in the continuous domain, we introduce a sampling heuristic that results in similar gains.
-Let $\bm{\epsilon}\sim\mathcal{N}(0,I)$ be a standard multivariate Gaussian random variable and let $\mathbf{x}=f(\bm{\epsilon})$, where $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a differentiable function.
-By the change of variables formula, the probability density of $\mathbf{x}$ is given by: $p(\mathbf{x})=p(\bm{\epsilon})\cdot\left|\det\left(\frac{\partial f(\bm{\epsilon})}{\partial\bm{\epsilon}}\right)\right|^{-1}$, where $\frac{\partial f(\bm{\epsilon})}{\partial\bm{\epsilon}}$ is the Jacobian matrix of $f$ at $\bm{\epsilon}$.
-In theory, to apply temperature sampling with temperature $\tau$ we should raise both terms to the power $1/\tau$, but the Jacobian term is untractable.
-In practice, we noticed that we can apply temperature only on the Gaussian noise sampling (which is equivalent to sampling from a Gaussian with variance $1/\sqrt{\tau}$) and get significant improvements in metrics.
-The effects of gaussian temperature are further discussed in Section 5.3.
-Table 1: Speech compression models. Our VAE is on par with the VQ-VAE on acoustic quality (MOSNeT) and outperforms it on semantic discriminability (ABX).
-Model Type Dims / RVQ Frame Rate (Hz) Bitrate (kbit/s) MOSNet ($\uparrow$) ABX ($\downarrow$)
-VQ-VAE (Mimi) 8 RVQ 12.5 1.1kbps 3.11 9.4%
-VAE 32 dims 12.5 - 3.15 8.1%
+To replicate this behavior in the continuous domain, we introduce a sampling heuristic that results in comparable gains.
+Similarly to the GAN noise truncation trick presented in  Brock et al.
+(2018), we sample more from the high probability zone of the Gaussian to trade diversity for fidelity.
+While the GAN truncation trick truncates the Gaussian noise such that values outside of a certain range are redrawn, we chose to reduce the variance of the Gaussian noise instead.
+This is mathematically equivalent to applying a temperature $\tau$ to the Gaussian if we change the standard deviation to $\sqrt{\tau}$.
+This makes temperature values between the discrete and continuous setups somewhat comparable, and we found that using a temperature of .8 for speech continuation was bringing good results in both setups.
+The effects of gaussian temperature are further discussed in Section C.
+4.6 Latent Classifier Free Guidance
+Classifier Free Guidance (CFG) (Ho and Salimans, 2022) is known to improve the generation quality of conditioned generative models.
+It can be applied for diffusion and flow matching models on the sampling trajectory as well as on the logits of autoregressive language models (Kreuk et al., 2023).
+Since CFG cannot be applied on the trajectory of 1-step consistency models we decide to apply the CFG on the outputs of the Backbone and Short-Context transformers.
+Formally, given $\mathbf{C}$ a conditioning and $\alpha$ the CFG coefficient, we compute for every $s$ of the sequence $\mathbf{Z}^{s}_{\text{CFG}}=\mathbf{Z}^{s}_{\emptyset}+\alpha(\mathbf{Z}^{s}_{C}-\mathbf{Z}^{s}_{\emptyset})$ and then we generate $\hat{\mathbf{x}}^{s}$ with the consistency head conditioned on $\mathbf{Z}^{s}_{\text{CFG}}$.
+We call this method Latent CFG, as it operates on the latent variable $\mathbf{Z}^{s}$ instead of the model output.
+It has been introduced in the video-to-audio model SoundReactor (Saito et al., 2025).
+4.7 Latent Distillation
+Once a teacher model has been trained and the desired classifier free guidance (CFG) coefficient has been selected for inference, we distill the CFG-guided teacher into a student model to avoid the need to double the batch size during inference when using CFG.
+To this end, we distill only the backbone transformer and directly copy the teacher’s MLP head into the student.
+The distillation objective for the student backbone is an $\ell_{2}$ loss between the latent representation $\mathbf{Z}^{s}_{\text{distill}}$ produced by the student backbone and the CFG-guided latent representation $\mathbf{Z}^{s}_{\text{CFG}}$ produced by the teacher.
+Additionally, the student backbone transformer may contain fewer layers than the teacher.
+In practice for Pocket TTS, we distill a text-to-speech model with 24 transformer layers into a student model with only 6 layers, using a latent CFG coefficient of $\alpha=1.5$.
+See Sec.
+F for more details.
 5 Experiments and results
-We now present experimental results on both speech and music continuation.
-We generate 27 seconds following a prompt of 3 seconds.
-All detailed hyperparameters are in the Appendix Sec.
-C.
-All metrics reported with confidence intervals use 95% confidence.
-Metrics without confidence intervals are metrics computed over the entire test dataset.
 5.1 Speech Continuation
@@
 Unlike Mimi, which applies this loss only to the first codebook, we extend it to the entire latent representation.
+Table 1: Speech compression models.
+Our VAE is on par with the VQ-VAE on acoustic quality (MOSNeT) and outperforms it on semantic discriminability (ABX), PESQ Rix et al.
+(2001) and STOI Taal et al.
+(2011) and a MUSHRA for acoustic quality.
+Model Type Dims / RVQ Frame Rate (Hz) Bitrate (kbit/s) MOSNet ($\uparrow$) ABX ($\downarrow$) PESQ ($\uparrow$) STOI ($\uparrow$) Acoustic Qual. ($\uparrow$)
+VQ-VAE (Mimi) 8 RVQ 12.5 1.1kbps 3.11 9.4% 2.13 0.87 57.7 $\pm$ 1.3
+VAE 32 dims 12.5 - 3.15 8.1% 2.42 0.90 66.0 $\pm$ 1.4
 Model and dataset: Starting from Helium-1 (Kyutai, 2025), a pretrained 2B parameters multilingual text LM as backbone, we train on French and English speech data following Défossez et al.
@@
 For speech generation, we didn’t notice any gains from introducing a short context transformer and noising the latents before feeding them to the backbone, resulting in a simpler model architecture.
+Table 2: Comparison of speech continuation models: 8-RVQ RQ-transformer vs 1-step Consistency model head, with 2 temperature options.
+Model Type Sampling Overall Sampler % Time in PPX ($\downarrow$) VERT ($\downarrow$) Acoustic Meaningfulness
+temperature Speedup ($\uparrow$) Speedup ($\uparrow$) Sampler ($\downarrow$) Quality($\uparrow$) Elo ($\uparrow$) Rank ($\downarrow$)
+Reference $20.2$ $25.2$ $4.02\pm 0.11$ $2180\pm 30$
+RQ-transformer 8 RVQ 1.0 $\times 1.0$ $\times 1.0$ 26.7% $52.4$ $36.3$ $2.42\pm 0.12$ $1841\pm 25$ 4
+RQ-transformer 8 RVQ 0.8 $\times 1.0$ $\times 1.0$ 26.7% $26.8$ $33.1$ $2.75\pm 0.14$ $1870\pm 30$ 3
+CALM - Consistency - 1 step 1.0 $\times 1.3$ $\times 12.3$ 2.9% $42.9$ $34.3$ $2.82\pm 0.13$ $1947\pm 28$ 2
+CALM - Consistency - 1 step 0.8 $\times 1.3$ $\times 12.3$ 2.9% $\bm{23.8}$ $\bm{31.2}$ $\bm{3.45}\pm\bm{0.14}$ $\bm{2023}\pm\bm{27}$ $\bm{1}$
 Results: Tab.
@@
 2 shows that the 1-step Consistency model outperforms the RQ-Transformer with 8 RVQ on all our automatic and human based metrics as well as on speed.
-For automatic metrics, we compute PPX and VERT Lakhotia et al.
+For automatic metrics, we compute PPX and VERT as introduced by Lakhotia et al.
 (2021).
 The PPX metric measures the semantic meaningfulness of the generated speech.
-To do so we generate 1000 excerpts of speech of 30 second, we use Whisper Radford et al.
-(2022) to compute textual transcriptions and finally compute the negative log-likelihood of the text tokens with a Mistral 7B LLM Jiang et al.
+To do so we generate 1000 excerpts of speech of 30 second, we use Whisper (Radford et al., 2022) to compute textual transcriptions and finally compute the negative log-likelihood of the text tokens with a Mistral 7B LLM Jiang et al.
 (2023) and convert it to Perplexity.
-Because a model that generates poorly diverse but good quality sentences would perform well on the PPX metric, the authors of Lakhotia et al.
-(2021) introduce the VERT metric (for diVERsiTy) which is a geometric mean of self- and auto-BLEU metrics.
+Because a model that generates poorly diverse but good quality sentences would perform well on the PPX metric, the authors of (Lakhotia et al., 2021) introduce the VERT metric (for diVERsiTy) which is a geometric mean of self- and auto-BLEU metrics.
 We use the official implementation from the fairseq Ott et al.
@@
 These pairwise preferences are used to compute an Elo score (see Sec.
-B).
+G).
 Notably, we note a clear quality and meaningfulness improvement with our temperature method.
@@
 The overall gain to perform a full inference of 30 seconds is $\times 1.3$.
-Table 2: Comparison of speech continuation models: 8-RVQ RQ-transformer vs 1-step Consistency model head, with 2 temperature options.
-Model Type Sampling Overall Sampler % Time in PPX ($\downarrow$) VERT ($\downarrow$) Acoustic Meaningfulness
-temperature Speedup ($\uparrow$) Speedup ($\uparrow$) Sampler ($\downarrow$) Quality($\uparrow$) Elo ($\uparrow$) Rank ($\downarrow$)
-Reference $20.2\pm 0.1$ $25.2\pm 0.1$ $4.02\pm 0.11$ $2180\pm 30$
-RQ-transformer 8 RVQ 1.0 $\times 1.0$ $\times 1.0$ 26.7% $52.4\pm 0.2$ $36.3\pm 0.2$ $2.42\pm 0.12$ $1841\pm 25$ 4
-RQ-transformer 8 RVQ 0.8 $\times 1.0$ $\times 1.0$ 26.7% $26.8\pm 0.1$ $33.1\pm 0.1$ $2.75\pm 0.14$ $1870\pm 30$ 3
-CALM - Consistency - 1 step 1.0 $\times 1.3$ $\times 12.3$ 2.9% $42.9\pm 0.2$ $34.3\pm 0.2$ $2.82\pm 0.13$ $1947\pm 28$ 2
-CALM - Consistency - 1 step 0.8 $\times 1.3$ $\times 12.3$ 2.9% $\bm{23.8}\pm\bm{0.1}$ $\bm{31.2}\pm\bm{0.1}$ $\bm{3.45}\pm\bm{0.14}$ $\bm{2023}\pm\bm{27}$ $\bm{1}$
+Temperature Sampling and speaker similarity: In Sec.
+C, we show that our gaussian temperature sampling heuristic has similar effects on speaker similarity than the temperature sampling of the discrete model.
+5.2 Text-to-Speech (TTS)
+VAE, model and dataset: We use the same VAE as for speech continuation.
+Our TTS CALM builds on a 300M backbone transformer and uses the same architecture for the 10M parameters consistency sampling head.
+The text is fed to the backbone as a prefix with SentencePiece model (Kudo and Richardson, 2018) with a vocabulary size of 4k.
+The training data is a mix of public datasets totalizing to 88k hours of speech that is detailed in Sec.
+D.
+Results: We evaluate on the Librispeech test-clean set using the same protocol as F5-TTS (Chen et al., 2025).
+We compare against four baselines: F5-TTS (Chen et al., 2025), DSM (Zeghidour et al., 2025), DiTAR (Jia et al., 2025), and SALAD (Turetzky et al., 2024).
+For DiTAR and SALAD, we report the paper results since the models are closed-source.
+We report Word Error Rate (WER) and Character Error Rate (CER) using Whisper-large-v3 (Radford et al., 2022), Speaker Similarity using WavLM-large (Chen et al., 2021b) as well as the results of a MUSHRA test for acoustic quality and a pairwise audio test for speaker similarity.
+Table 3: Text-to-Speech models. Our CALM model with 1-step LSD outperforms baselines on WER, CER and Acoustic Quality. Results for Pocket TTS are in Sec. F.
+Model Num. parameters WER CER SIM Acoustic Speaker SIM
+($\downarrow$) ($\downarrow$) ($\uparrow$) Quality ($\uparrow$) Human ELO ($\uparrow$)
+Reference 2.23 0.69 61.8 $\pm$ 2.4 1953 $\pm$ 24
+Reference (with VAE) 0.57
+F5 TTS (NFE=32) (Chen et al., 2025) 336M 2.42 0.66 54.7 $\pm$ 2.8 2032 $\pm$ 18
+DSM (16 RVQ CFG=3 750M 1.95 0.67 60.2 $\pm$ 2.4 2112 $\pm$ 20
+w.r.t text and audio prompt) (Zeghidour et al., 2025)
+DiTAR (NFE=10) (Jia et al., 2025) 600M 2.39 0.67
+SALAD (NFE=20) (Turetzky et al., 2024) 350M 0.74 0.54
+CALM w/ LSD (NFE=1, CFG=1.5 w.r.t text) 313M 1.81 0.57 0.52 61.1 $\pm$ 2.3 1966 $\pm$ 23
+Our CALM model with 1-step LSD (Boffi et al., 2025) outperforms baselines on WER, CER and Acoustic Quality but obtains a low speaker similarity score.
+This can be partially explained by the fact that when computing the speaker similarity between the reference prompt and the reference utterance that goes through the VAE (the second line of the tab) we obtain a similarity of 0.57.
+Yet, we demonstrate in Tab.
+1 that our VAE faithfully reconstructs speech.
+Due to this surprising behavior, we decide to measure speaker similarity with a human study.
+We observe that all measured methods beat the ground truth, which means that they preserve well the voice of the audio prompt.
 5.3 Music Continuation
 Dataset: We use a randomly selected subset of 400K songs (approximately 20K hours with 32kHz mono format) from the LAION-Disco-12M dataset, ensuring broad coverage across musical genres.
-VAE: Our variational autoencoder (VAE) and codec architecture are adapted from the Mimi codec (Défossez et al., 2024b), originally designed for 24kHz speech at 12.5Hz.
+VAE: Our variational autoencoder (VAE) and codec architecture is adapted from the Mimi codec (Défossez et al., 2024b), originally designed for 24kHz speech at 12.5Hz.
 We trained it to compress 32kHz mono music with a 25Hz frame rate.
-We experiment with bottleneck sizes of 96 and 128 dimensions.
-For comparison, MusicGen’s EnCodec model (Copet et al., 2023) also operates at 32kHz but uses a 4-level RVQ at 50Hz.
-In Tab.
-3, we report reconstruction metrics (audio ViSQOL Chinen et al.
-(2020) and SISNR), showing that a 32-dim VAE matches MusicGen’s codec, and that at least 96 dimensions are needed for our VAE to match the quality of the 32-level RVQ configuration.
-Table 3: Music compression models. At least 96 VAE latent dimensions are required to outperform the 32-RVQ codec on reconstruction metrics. EnCodec has been retrained on our dataset.
-Model Type Dims / RVQ Frame Rate (Hz) Bitrate (kbps) VisQOL ($\uparrow$) SISNR ($\uparrow$)
-EnCodec Copet et al. (2023) 4 RVQ 50 2.2 2.41 5.62
-VQ-VAE (inspired from Mimi) 32 RVQ 25 8.8 3.63 9.61
-VAE 32 dims 25 2.23 5.51
-VAE 96 dims 25 3.65 9.76
-VAE 128 dims 25 4.01 10.3
+Details and metrics are described in Sec.
+B.
 Table 4: Comparison of music generation models across speed and quality metrics for 30 seconds generation.
@@
 Model: Our music CALM builds on the MusicGen Medium backbone, a 1.35B parameter Transformer (see Sec.
-C for all the hyperparameters).
+H for all the hyperparameters).
 We compare our method against: (1) Two discrete models using 32 and 16 RVQ codecs, each employing the same 1.35B backbone and a RQ-transformer for parallel codes prediction; (2) a MusicGen Medium variant trained with EnCodec without textual conditioning, using the same backbone and a delay pattern for codebook interleaving.
@@
 The Enjoyment metric is an Elo score (see Sec.
-B), computed by making human rater choose their favorite music out of generated pairs with the same 3s prompt.
-The evaluation protocol uses the same amount of annotators and examples as the speech study.
+G), computed by making human rater choose their favorite music out of generated pairs with the same 3s prompt.
 We observe that CALM with consistency outperforms the 32 RVQ RQ-Transformer baseline on computed and human metrics while being $\times 1.9$ to $\times 2.2$ times faster for the overall speedup.
@@
 As well, we train a CALM model with a TrigFlow head instead of consistency and it outperforms all the models but to the price of a slow inference.
-Figure 2: Effect of the head batch multiplier value.
-Training a model (music consistency CALM) with a higher batch size multiplier fastens the convergence for the FAD metric.
-All the evaluations are done with 4 steps of consistency at inference time.
 An ablation study (Tab.
@@
 In Fig.
-2, we show that the FAD decreases much faster over time when we train consistency CALM models with a bigger head batch multiplier.
+3, we show that the FAD decreases much faster over time when we train consistency CALM models with a bigger head batch multiplier.
 All evaluations are done with 4 steps of consistency.
 We keep the value of 8 for all of our experiments as a higher value would lead to out of memory issues at training time.
-Scalability: We show in Sec. A.2 that CALM does improve with a bigger 3B parameters backbone. However, we leave a complete scalability study for future work.
-Necessity of Consistency for a fast inference: We show in Tab.
-6 that the consistency framework largely outperforms the TrigFlow framework for 10 inference steps and less (the regime where the Real time factor is greater than 1).
-Table 6: Generation efficiency of TrigFlow CALM and Consistency CALM models.
-We compute the inference time to generate 30 seconds of audio, the corresponding Real Time Factor (RTF) as well as the FAD metric for different numbers of inference steps.
-For less than 10 steps (i.e.
-when streaming applications are possible, the consistency method outperforms the TrigFlow method).
-# Steps Time (s) RTF FAD (TrigFlow) FAD (Consistency)
-1 16.7 $\times 1.80$ $0.83\pm 0.04$
-4 20.4 $\times 1.47$ $28.83\pm 0.20$ $0.71\pm 0.05$
-10 27.7 $\times 1.08$ $4.62\pm 0.07$ $0.73\pm 0.05$
-25 46.1 $\times 0.65$ $0.79\pm 0.04$ $0.96\pm 0.06$
-50 76.8 $\times 0.39$ $0.74\pm 0.05$ $1.46\pm 0.06$
-100 136.4 $\times 0.22$ $0.64\pm 0.04$ $2.05\pm 0.07$
-5.3 Effect of gaussian temperature sampling
+Necessity of Consistency for a fast inference: We show in Sec.
+E.3 that the consistency framework largely outperforms the TrigFlow framework for 10 inference steps and less (the regime where the Real time factor is smaller than 1).
+Scalability: We show in Sec. E.4 that CALM does improve with a bigger 3B parameters backbone. However, we leave a complete scalability study for future work.
+5.4 Text-to-music generation
+We compare CALM on the task of text-to-music generation.
+Since our dataset do not have any text labeling, we use CLAP (Elizalde et al., 2023) as a prefix conditioning and train MusicGen, our RQ-Transformer 32 RVQ baseline as well as CALM.
+During training, we drop the conditioning $20\%$ of the time in order to perform Classifier Free Guidance (Ho and Salimans, 2022) at inference.
+We use CLAP in audio mode when training.
+We use the same training data, VAE/codecs and hyperparameters as the models reported in Tab.
+4.
+In Tab.
+6, we show the results of our CLAP conditioned model evaluated on our internal test set with the CLAP conditioning being used in audio mode.
+As well, we report the results of our model and several baselines on the text-to-music benchmark on the MusicCaps dataset Agostinelli et al.
+(2023) in Sec.E.6.
+Table 6: Text-to-music generation models on our test set where CLAP is used in audio mode.
+Model Overall Sampler % time in FAD ($\downarrow$) Acoustic Enjoyment
+Speedup ($\uparrow$) Speedup ($\uparrow$) Sampler ($\downarrow$) Quality ($\uparrow$) Elo ($\uparrow$) Rank ($\downarrow$)
+Reference 3.75 $\pm$ 0.15 2104 $\pm$ 24
+RQ-transformer 32 RVQ (CFG=3) $\times$ 1.0 $\times$ 1.0 57.7% 0.93 $\pm$ 0.07 3.16 $\pm$ 0.15 1998 $\pm$ 24 1
+CALM - Consistency - 4 Steps (CFG=2) $\times$ 1.9 $\times$ 5.4 20.1% 0.91 $\pm$ 0.08 3.11 $\pm$ 0.14 1998 $\pm$ 24 1
+MusicGen Medium $\times$ 1.3 0.0% 1.93 $\pm$ 0.14 2.54 $\pm$ 0.14 1946 $\pm$ 25 3
+6 Conclusion
+We present Continuous Audio Language Models (CALM), a novel framework for autoregressive audio generation that operates directly in the continuous latent space of a VAE, bypassing the limitations of discrete quantization.
+Replacing RVQ or diffusion heads with consistency models significantly reduces inference cost while improving sample quality as shown by our experiments.
+Our architecture combines noise-injected long-term context and clean short-term context, implemented via a dual-transformer design.
+We introduce practical innovations such as temperature sampling and a head batch multiplier to further improve sampling quality and training efficiency.
+We demonstrate the effectiveness of our approach across both speech and music generation tasks.
+Our results suggest that continuous modeling offers a compelling alternative to discrete tokenization for high-quality, efficient, and scalable autoregressive audio generation.
+7 Reproducibility statement
+We will release the training code in order to retrain CALM. As well, we provide all the model hyperparameters in Sec. H.
+Appendix A Lagrangian Self-Distillation
+Lagrangian Self-Distillation (LSD) (Boffi et al., 2025) extends the consistency-model framework by introducing an additional time parameter $s$, enabling the model to learn mappings between arbitrary points along the probability flow trajectory rather than only from noisy inputs to clean data.
+An LSD model is a neural network $f_{\phi}(\mathbf{x},t,s)$ that predicts the state of the PF-ODE solution at time $s$ given its state $\mathbf{x}_{t}$ at time $t$: $\forall(s,t)\in[0,1]^{2},f_{\phi}(\mathbf{x}_{t},t,s)=\mathbf{x_{s}}$.
+In particular, with $t=1,s=0$ and by sampling $\mathbf{x}_{1}\sim\mathcal{N}(0,I)$ we can sample from the data distribution with $f_{\phi}(\mathbf{x}_{1},t=1,s=0)$.
+In (Boffi et al., 2025), the authors define $\mathbf{F}_{\phi}(\mathbf{x},t,s)$ as
+$f_{\phi}(\mathbf{x},t,s)=\mathbf{x}+(s-t)F_{\phi}(\mathbf{x},t,s).$ (4)
+To train a LSD model from scratch we need to combine a flow matching loss as well as a LSD loss: $\mathcal{L}=\mathcal{L}_{FM}+\mathcal{L}_{LSD}$.
+With an adaptive weighting loss $w_{\psi}(t,s)$ and our notations they derive as:
+$\mathcal{L}_{\text{FM}}(\phi,\psi)=\mathbb{E}_{\mathbf{x}_{0}\sim p_{\text{data}},\,\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),\,t\sim\mathcal{U}(0,1)}\left[e^{-w_{\psi}(t,t)}\left\|F_{\phi}(\mathbf{x}_{t},t,t)-\left(\alpha^{\prime}_{t}\mathbf{x}_{0}+\sigma^{\prime}_{t}\bm{\epsilon}\right)\right\|_{2}^{2}+w_{\psi}(t,t)\right].$ (5)
+and
+$\mathcal{L}_{\text{LSD}}(\phi,\psi)=\mathbb{E}_{\mathbf{x}_{0},\bm{\epsilon},t,s}\left[e^{-w_{\psi}(t,s)}\left\|\partial_{s}f_{\phi}(\mathbf{x}_{t},t,s)-F_{\phi^{-}}(f_{\phi}(\mathbf{x}_{t},t,s),s,s)\right\|_{2}^{2}+w_{\psi}(t,s)\right].$ (6)
+where $\mathbf{x}_{0}\sim p_{\text{data}},\,\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),t\sim\mathcal{U}(0,1),s\sim\mathcal{U}(0,t)$.
+In practice, we compute the flow matching loss on 75% of a batch and the LSD loss on the 25% left.
+Appendix B Our Music VAE
+Table 7: Music compression models. At least 96 VAE latent dimensions are required to outperform the 32-RVQ codec on reconstruction metrics. EnCodec has been retrained on our dataset.
+Model Type Dims / RVQ Frame Rate (Hz) Bitrate (kbps) VisQOL ($\uparrow$) SISNR ($\uparrow$)
+EnCodec Copet et al. (2023) 4 RVQ 50 2.2 2.41 5.62
+VQ-VAE (inspired from Mimi) 32 RVQ 25 8.8 3.63 9.61
+VAE 32 dims 25 2.23 5.51
+VAE 96 dims 25 3.65 9.76
+VAE 128 dims 25 4.01 10.3
+Our variational autoencoder (VAE) and codec architecture is adapted from the Mimi codec (Défossez et al., 2024b), originally designed for 24kHz speech at 12.5Hz.
+We trained it to compress 32kHz mono music with a 25Hz frame rate.
+We experiment with bottleneck sizes of 96 and 128 dimensions.
+For comparison, MusicGen’s EnCodec model (Copet et al., 2023) also operates at 32kHz but uses a 4-level RVQ at 50Hz.
+In Tab.
+7, we report reconstruction metrics (audio ViSQOL Chinen et al.
+(2020) and SISNR), showing that a 32-dim VAE matches MusicGen’s codec, and that at least 96 dimensions are needed for our VAE to match the quality of the 32-level RVQ configuration.
+Appendix C Effect of gaussian temperature sampling
 We evaluate how our proposed gaussian temperature sampling method affects acoustic diversity for consistency models, in comparison with temperature sampling on its discrete counterpart.
-More precisely, we compute WavLM speaker embeddings (Chen et al., 2021) over 100 unprompted speech generations with different values of temperature, both for an RQ-transformer and CALM.
+More precisely, we compute WavLM speaker embeddings (Chen et al., 2021b) over 100 unprompted speech generations with different values of temperature, both for an RQ-transformer and CALM.
 In both cases, the inner monologue text stream is generated with 0.8 temperature to ensure text diversity.
@@
 less pairwise speaker similarity) as temperature increases.
-6 Conclusion
-We present Continuous Audio Language Models (CALM), a novel framework for autoregressive audio generation that operates directly in the continuous latent space of a VAE, bypassing the limitations of discrete quantization.
-Replacing RVQ or diffusion heads with consistency models significantly reduces inference cost while improving sample quality as shown by our experiments.
-Our architecture combines noise-injected long-term context and clean short-term context, implemented via a dual-transformer design.
-We introduce practical innovations such as temperature sampling and a head batch multiplier to further improve sampling quality and training efficiency.
-We demonstrate the effectiveness of our approach across both speech and music generation tasks.
-Our results suggest that continuous modeling offers a compelling alternative to discrete tokenization for high-quality, efficient, and scalable autoregressive audio generation.
-Appendix A Supplementary Experiments
-A.1 Short-context transformer window
-To study the influence of the context window $K$ of the short-context transformer, we performed an hyperparameter search with $K=5,10,20,40$ and trained our models for 500K training steps.
-We observe that finding an optimal value of $K$ is not critical even though it seems that 5 is not enough.
+Appendix D Data used for the text-to-speech model
+The dataset used to train our TTS models is composed of AMI Carletta (2007), EARNINGS22 Del Rio et al.
+(2022), GIGASpeech Chen et al.
+(2021a), SPGISpeech O’Neill et al.
+(2021), TED-LIUM Hernandez et al.
+(2018), VoxPopuli Wang et al.
+(2021), LibriHeavy Beck et al.
+(2024), and Emilia Author and Contributors (2023).
+It results into 88k hours of audio.
+Appendix E Supplementary Experiments
+E.1 Head batch multiplier value
+Figure 3: Effect of the head batch multiplier value.
+Training a model (music consistency CALM) with a higher batch size multiplier fastens the convergence for the FAD metric.
+All the evaluations are done with 4 steps of consistency at inference time.
+E.2 Short-context transformer window
+To study the influence of the context window $K$ of the short-context transformer, we performed an hyperparameter search with $K=5,10,20,40$ and trained our models for 500K training steps (see Tab.
+8).
+The model is the music continuation CALM and all evaluations are done with 4-steps of consistency.
+We observe that finding an optimal value of $K$ is not critical even though it seems that 5 is too small enough.
 Table 8: FAD after 500K training steps for different short-context Transformer contexts $K$.
@@
 40 $0.78\pm 0.04$
-A.2 Scalability of CALM
+E.3 Comparison of TrigFlow and Consistency on inference speed and quality criteria
+In Tab.
+9 we compare the inference speed for both TrigFlow CALM and Consistency CALM on the task of music continuation on our test set.
+We see that below the 10 steps regime, only the Consistency model performs well.
+Moreover, around 10 steps is the limit for streaming applications where the real time factor (RTF) is smaller than 1.
+This justifies the need of consistency modeling (instead of diffusion or flow matching) for good quality streaming generation.
+Table 9: Generation efficiency of TrigFlow CALM and Consistency CALM models.
+We compute the inference time to generate 30 seconds of audio, the corresponding Real Time Factor (RTF) as well as the FAD metric for different numbers of inference steps.
+For streaming (RTF<1), only consistency generates good quality audio.
+# Steps Time (s) RTF FAD (TrigFlow) FAD (Consistency)
+1 16.7 $0.56$ $0.83\pm 0.04$
+4 20.4 $0.68$ $28.83\pm 0.20$ $0.71\pm 0.05$
+10 27.7 $0.92$ $4.62\pm 0.07$ $0.73\pm 0.05$
+25 46.1 $1.54$ $0.79\pm 0.04$ $0.96\pm 0.06$
+50 76.8 $2.56$ $0.74\pm 0.05$ $1.46\pm 0.06$
+100 136.4 $4.55$ $0.64\pm 0.04$ $2.05\pm 0.07$
+E.4 Scalability of CALM
 In order to see if all the hyperparameters of our CALM method transfer well to more model parameters, we trained a consistency music CALM model with a larger 3B backbone (with a model dimension of 2048, 32 heads and 48 layers) as well as a discrete RQ-Transformer with 32-RVQ model with the same backbone size.
 For the CALM model, we use 4 consistency steps at inference.
-The results are as follow (we compare with the main model that has a 1.3B backbone):
+Tab.
+10 shows that the FAD are improving with a bigger backbone (3B vs 1.3B) in similar proportions for both discrete-based (RQ-Transformer) and continuous-based (CALM) models.
 Table 10: Scalability of the backbone for CALM and RQ-Transformer methods.
@@
 32 RVQ RQ-Transformer with 1.3B backbone $1.06\pm 0.06$
-Appendix B Human evaluation methods
+E.5 Text-to-Speech ablation study
+Model Num. parameters WER CER SIM MUSHRA Speaker SIM
+($\downarrow$) ($\downarrow$) ($\uparrow$) ($\uparrow$) (Human eval ELO ($\uparrow$)
+CALM w/ LSD (NFE=1, CFG=1.5 w.r.t text) 313M 1.81 0.57 0.52 61.1 $\pm$ 2.3 1966 $\pm$ 23
+CALM w/ LSD (NFE=1, no CFG) 313M 2.39 0.90 0.52 55.5 $\pm$ 2.2 1991 $\pm$ 26
+CALM w/ LSD (NFE=1, CFG=1.25 w.r.t text and audio prefix) 313M 1.86 0.59 0.54 56.2 $\pm$ 2.2 1995 $\pm$ 27
+CALM w/ Consistency (NFE=1, CFG=1.5 w.r.t text) 313M 1.93 0.62 0.40 46.8 $\pm$ 2.5 1900 $\pm$ 30
+We can see in Tab.
+10 that at this smaller scale ($\sim$300M), Lagrangian Self-Distillation models provide better audio quality than Consistency models.
+We also conclude that the latent CFG has a significant impact over the WER.
+As well, applying the latent CFG on the audio prefix improves the speaker similarity but at the cost of audio quality.
+All models were run with sampling from a Gaussian with 0.7 temperature (i.e., multiplying the standard Gaussian noise by $\sqrt{0.7}$).
+E.6 Text-to-Music evaluation on MusicCaps
+We report the results on the MusicCaps dataset Agostinelli et al.
+(2023) and report as well the results of MusicLM (Agostinelli et al., 2023), the original text-to-music MusicGen Medium (Copet et al., 2023), AudioLDM2 (Liu et al., 2024), Noise2Music (Huang et al., 2023), Jen-1 (Li et al., 2025) as well as MusicFlow (Prajwal et al., 2024) on the FAD, the KL Divergence (KLD) metric based on the pre-trained audio model PANN (Kong et al., 2019) as well as the CLAP cosine similarity that measures the matching between the generated music and text description.
+The results of Tab.
+11 show that even though our goal is not to specifically build the best text-to-music model, simply applying CALM on our dataset with a CLAP conditioning leads to competitive results.
+The metrics from the models that we did not train are reported from the MusicFlow (Prajwal et al., 2024) paper.
+Table 11: Text-to-music results on the MusicCaps dataset.
+Model FAD ($\downarrow$) KLD ($\downarrow$) CLAP ($\uparrow$)
+Reference 0.30
+MusicLM (Agostinelli et al., 2023) 4.00
+MusicGen Medium (Copet et al., 2023) 3.40 1.23 0.37
+AudioLDM2 (Liu et al., 2024) 3.13 1.20 0.43
+Noise2Music (Huang et al., 2023) 2.10
+Jen-1 (Li et al., 2025) 2.00 1.29
+MusicFlow (Unidirectional LM + FM) (Prajwal et al., 2024) 2.69 1.23 0.52
+MusicGen Medium (retrained with CLAP) 2.70 1.37 0.39
+RQ-transformer 32 RVQ 2.56 1.35 0.43
+CALM - Consistency - 4 Steps 2.14 1.30 0.44
+Appendix F Pocket TTS
+Given a 313M parameters text-to-speech CALM model that has a 24 layers backbone transformer (the teacher), we do latent distillation to a 6 layers transformer backbone (the student) and a CFG coefficient of $\alpha=1.5$ applied to the teacher.
+We keep the same MLP sampling head.
+The final model, named Pocket TTS has a size of 90M parameters while the VAE has 20M parameters.
+In Tab.12, we put 100M parameters to include the VAE decoder in the parameter count.
+We evaluate Pocket TTS on the Librispeech test-clean set following the same protocol as F5-TTS (Chen et al., 2025), with the difference that we cleaned the audio input using Adobe Enhance Speech [^1] https://podcast.adobe.com/en/enhance to obtain 24kHz high-quality audio.
+We evaluate all baselines with the enhanced samples [^2] https://huggingface.co/datasets/kyutai/librispeech_test_clean_enhanced.
+We compare against three baselines: F5-TTS (Chen et al., 2025), DSM (Zeghidour et al., 2025), Chatterbox Turbo [^3] https://www.resemble.ai/chatterbox-turbo/ and Kokoro TTS [^4] https://huggingface.co/hexgrad/Kokoro-82M.
+We report Word Error Rate (WER) using Whisper-large-v3(Radford et al., 2022), as well as the results of a pairwise human evaluation for audio quality and speaker similarity.
+For audio quality, we ask raters “Which of the two audio clips has the best audio quality?”, and for speaker similarity, we ask “Which of the two audio clips sounds more similar to the reference audio clip in terms of voice characteristics?” and provide the voice prompt as a reference.
+Table 12: Comparison of TTS models on Librispeech test-clean.
+Model Param size WER Audio Quality Speaker Sim Faster than
+(gen. only) ($\downarrow$) (ELO) ($\uparrow$) (ELO) ($\uparrow$) real-time CPU
+F5-TTS (Chen et al., 2025) 336M 2.21 $1949\pm 27$ $1946\pm 26$ $\times$
+DSM 750M 1.84 $1959\pm 25$ $2037\pm 21$ $\times$
+Chatterbox Turbo 350M 3.24 $2055\pm 23$ $2012\pm 22$ $\times$
+Kokoro 82M 1.93 No voice cloning $\checkmark$
+Pocket TTS (ours) 100M 1.84 $2016\pm 25$ $1898\pm 26$ $\checkmark$
+As seen in Tab.
+12, Pocket TTS has the lowest Word Error Rate, a better Audio Quality than the ground truth F5-TTS and DSM, as well as an on-par Speaker Similarity with the ground truth while being a significantly smaller model than competitors, and being the only one that can run faster than real-time on CPU (we tested on Apple M3 and Intel core ultra 7 165H).
+We invite the reader to check the blog post where the one-line installation of the model is provided: kyutai.org/pocket-tts-technical-report.
+Appendix G Human evaluation methods
 Audio clips are always 30s second total, with a 3s prompt coming from a ground truth audio.
@@
 Raters were payed £9 / hour for their contribution.
-B.1 Speech continuation
+G.1 Speech continuation
 Acoustic quality assessment: How would you rate the overall quality of this audio clip?
@@
 Elo scores are Bayesian estimates of the posterior mean in a Bradley-Terry model.
-B.2 Music continuation
+G.2 Music continuation and CLAP-to-music
 Acoustic quality assessment: How would you rate the overall quality of this music?
@@
 Elo scores are Bayesian estimates of the posterior mean in a Bradley-Terry model.
-B.3 Bayesian Elo Score
+G.3 Bayesian Elo Score
 The Meaningfulness metric for speech and the Enjoyment metric for music are both Bayesian Elo Scores.
@@
 $P(A>B)=\frac{S_{A}}{S_{A}+S_{B}}$ (8)
-which is a Bradley-Terry (Bradley & Terry, 1952) model.
+which is a Bradley-Terry (Bradley and Terry, 1952) model.
 There are a few different methods to estimate the parameters of a Bradley-Terry model.
-We use the iterative one from (Caron & Doucet, 2010) where $S_{A}^{0}$ follows a Gamma prior with parameters $\alpha^{0},\beta^{0}$.
+We use the iterative one from (Caron and Doucet, 2010) where $S_{A}^{0}$ follows a Gamma prior with parameters $\alpha^{0},\beta^{0}$.
 By denoting $w_{A}$ the number of times where method $A$ won against any other methods and $n_{AB}$ the number of times where $A$ and $B$ are compared, $S_{A}^{t}$ is computed with the following update rule until convergence:
@@
 Confidence interval are 95% confidence interval according to the posterior (the 2.5th and 97.5th percentiles).
-Appendix C Hyperparameters
+Appendix H Hyperparameters
 We present in Tab. 13 the parameters of the music and speech VAE used for CALMs. For CALMs and the RQ-Transformer based discrete LMs, the hyperparameters are presented in Tab. 14.
-Table 9: Model and training hyperparameters
-Music continuation Speech continuation
-Backbone Transformer
-Model dimension 1536 2560
-MLP dimension 6336 10560
-Number of heads 24 20
-Number of layers 48 24
-Learning rate $1\cdot 10^{-4}$ $5\cdot 10^{-5}$
-Number of parameters 1.35B 2.2B
-RQ-Transformer head (RVQ model)
-Model dimension 1024 1024
-MLP dimension 4096 4096
-Number of heads 16 16
-Number of layers 6 6
-Number of parameters 701M 701M
-Consistency sampler head (ours)
-Number of layers 12 6
-MLP dimension 3072 512
-Gating SiLU SiLU
-Number of parameters 601M 10M
-Short context transformer (ours)
-Model dimension 1536
-MLP dimension 6336
-Number of heads 24
-Number of layers 4
-Context 10
-Number of parameters 113M
-Audio embedding manipulations
-Center and normalize
-Noise before entering backbone
-Training parameters
-Head batch multiplier 8 8
-Optimizer AdamW $\beta_{1}=0.9,\beta_{2}=0.95$ AdamW $\beta_{1}=0.9,\beta_{2}=0.95$
-Batch size 48 144
-Audio sample length 30s 300s
-LR Schedule cosine cosine
-Learning rate $1\cdot 10^{-4}$ $2\cdot 10^{-4}$
-GPU used 16 H100 48 H100
-Number of training steps 500k 150k
-Initial checkpoint Helium-1 (Kyutai, 2025)
-Inner monologue
-Acoustic delay - 2 frames (160 ms)
 Table 13: VAE hyperparameters
-Music VAE Speech VAE
+Music VAE Speech continuation VAE
 General
@@
 Learning rate $8\cdot 10^{-4}$ $8\cdot 10^{-4}$
+Table 14: Model and training hyperparameters
+Music Speech continuation Text-to-Speech
+Backbone Transformer
+Model dimension 1536 2560 1024
+MLP dimension 6336 10560 4096
+Number of heads 24 20 16
+Number of layers 48 24 24
+Learning rate $1\cdot 10^{-4}$ $5\cdot 10^{-5}$ 1e-4
+Number of parameters 1.35B 2.2B 302M
+RQ-Transformer head (RVQ model)
+Model dimension 1024 1024
+MLP dimension 4096 4096
+Number of heads 16 16
+Number of layers 6 6
+Number of parameters 701M 701M
+Consistency sampler head (ours)
+Number of layers 12 6 6
+MLP dimension 3072 512 512
+Gating SiLU SiLU SiLU
+Number of parameters 601M 10M 10M
+Short context transformer (ours)
+Model dimension 1536
+MLP dimension 6336
+Number of heads 24
+Number of layers 4
+Context 10
+Number of parameters 113M
+Audio embedding manipulations
+Center and normalize
+Noise before entering backbone
+Training parameters
+Head batch multiplier 8 8 8
+Optimizer AdamW $\beta_{1}=0.9,\beta_{2}=0.95$ AdamW $\beta_{1}=0.9,\beta_{2}=0.95$ AdamW $\beta_{1}=0.9,\beta_{2}=0.95$
+Batch size 48 144 128
+Audio sample length 30s 300s 60s
+LR Schedule cosine cosine cosine
+Learning rate $1\cdot 10^{-4}$ $2\cdot 10^{-4}$ 1e-4
+GPU used 16 H100 48 H100 8 H100
+Number of training steps 500k 150k 400k
+Initial checkpoint Helium-1 (Kyutai, 2025)
+Inner monologue
+Acoustic delay - 2 frames (160 ms) -